### v1

# Schrödinger Symmetry in Spherically-symmetric Static Mini-superspaces with Matter Fields

Abstract
Schrödinger symmetry emerged in a “fluid limit” from a full superspace to several mini-superspace models. We consider two spherically-symmetric static mini-superspace models with matter fields and verify the robustness of this emergent symmetry at the classical level: (i) Maxwell field with cosmological constant and (ii) $n$ massless scalar fields. We develop a method based on canonical transformations and show that: for model (i), 3D Schrödinger symmetry emerges, and the solution is the (anti-)de Sitter Reissner-Nordström spacetime, and for model (ii), (2+n)(2+n)D Schrödinger symmetry appears, and the solution is a generalized Janis–Newman–Winicour spacetime and its “interior”, a Kantowski-Sachs type closed universe. In the matter decoupling limit, both cases lead to 2D Schrödinger symmetry in different lapse functions and mini-superspace coordinates, which implies the covariance of Schrödinger symmetry. Finally, we propose a physical interpretation of the symmetry under Hamiltonian constraints $H$ and explain it with examples: Symmetry generators commuting with $H$ map a solution to another one, while those non-commuting with $H$ generate a new theory with the Schrödinger symmetry and the transformed configuration is a solution to the new theory. These support the robustness of the emergence of Schrödinger symmetry and open new possibilities for exploring quantum dynamics of matter and gravity based on the symmetry.
I Introduction
Symmetry has played a significant role in physics. In the quest for a theory of quantum gravity, where the fundamental principles are still unknown, symmetry should serve as a guiding principle. The characteristic symmetry of gravity is general covariance, the gauge symmetry of spacetime, which leads to the background independence and dynamics of gravity. Here, the Hamiltonian constraint $\mathcal{H}=0$ is the master key; it generates diffeomorphism Dirac , the dynamics of gravity ultimately reduce to it Peres1962 ; DeWitt:1967yk ; osti_4124259 ; Gerlach:1969ph ; Moncrief:1972cx , and it plays a central role in various crucial problems, such as the so-called “time problem” Kuchar:1991qf ; Isham:1992ms ; Hoehn:2020epv and the quantization of gravity DeWitt:1967yk ; osti_4124259 . Thus, the Hamiltonian constraint should be the essential ingredient in exploring the symmetry of (quantum) gravity.
Then, what is the physical symmetry in the configuration space of gravity? That is, what is the dynamical symmetry (not gauge symmetry), such as transferring a physical configuration to another? One idea is corner symmetry; the existence of a spacetime boundary transforms the gauge symmetries into physical ones, and charges like energy evaluated on that boundary form the symmetry algebra Ciambelli:2022cfr ; Freidel:2023bnj . The crucial point here is that, the vanishing of the bulk contribution is a consequence of $\mathcal{H}=0$. Therefore, exploring the physical symmetries associated with $\mathcal{H}=0$ from the bulk perspective is highly intriguing. The present paper studies this and focuses particularly on symmetries that emerge in a certain limit.
One framework to describe the configuration space of gravitational field and matter fields is superspace. The Hamiltonian constraint $\mathcal{H}=0$ as the dynamical equation in general relativity is the Hamilton-Jacobi equation that determines a trajectory in the configuration space (superspace) of geometries and matter fields on a hypersurface in spacetime Peres1962 ; DeWitt:1967yk ; osti_4124259 ; Gerlach:1969ph . Superspace is an abstract manifold with infinite dimension, determined by the constraint $\mathcal{H}$.
When geometric symmetries such as homogeneity or spherical symmetry are imposed on spacetime, the number of configuration modes becomes finite, resulting in a finite-dimensional superspace (mini-superspace). This corresponds to restricting the full superspace to collective modes or taking a fluid limit, and thus universal structures independent of the details of the system can be expected to emerge Oriti:2024elx ; BenAchour:2024gir . Furthermore, typical examples of mini-superspace models such as black holes and homogeneous universe exhibit thermodynamic properties. This suggests that underlying microscopic degrees of freedom may exist while their statistical behavior is governed by universal mechanisms.
In fact, Schrödinger symmetry appears in both the homogeneous isotropic mini-superspace and the spherically-symmetric static vacuum mini-superspace PhysRevLett.90.031302 ; Achour2019 ; Achour2020 ; Achour_2021 ; 1021468SciPostPhys101022 ; Achour:2021dtj ; BenAchour:2023dgj ; Christodoulakis:2013xha ; BenAchour:2022fif ; Geiller:2022baq . It is a conformal extension of Galilean invariance Niederer:1972zz , and the generators consist of momenta $p_{i}$, angular momenta $j_{ij}$, Galilean boost $b^{i}$, Hamiltonian $H=q_{+}$, dilatation $q_{0}$, and special conformal transformation $q_{-}$ (see Sec.II or Ref.Duval:2024eod for a review). It is known not only as the maximal symmetry group of the free Schrödinger equation Niederer:1972zz but also as a symmetry inherent in the Gross-Pitaevskii equation, which is the effective equation of quantum liquids Kolomeisky:2000zz . Note that the symmetry algebra contains the Hamiltonian constraint $q_{+}=H(=0)$ but is not a constraint algebra like the hypersurface deformation algebra Dirac . Interpreting this symmetry as a universal symmetry of quantum gravity in the fluid limit, we can study models and quantization based on the symmetry; the symmetry can be useful for restricting the type of interactions in effective models of quantum gravity BenAchour:2023dgj or for controlling the ambiguity that arises when quantizing mini-superspace models to resolve singularities Sartini:2021ktb .
Therefore, it is crucial to verify the robustness of the emergence of the Schrödinger symmetry from various perspectives. In particular, it is essential to check whether the symmetry is preserved when interaction with matter is considered. In this paper, as a first step, we consider spherically-symmetric static mini-superspace models with matter fields  [^1] Spherically-symmetric static mini-superspace models can describe black holes. In quantum theory, matter fields play non-trivial roles. Hawking radiation originates from quantum fluctuations of matter fields, and interaction between matter fields and gravity should be considered in order to understand the time evolution of the information of a collapsing matter during the evaporation Kawai:2015uya . The origin of the entropy-area law, which should be consistent with the mechanism of information recovery, is also likely to be relevant to effects of matter Yokokura:2022phq . Furthermore, high curvatures, say, around the central region, induce large quantum fluctuations of matter fields, generating a large backreaction to the geometry Kawai:2020rmt . Thus, incorporating matter fields is crucial not only for examining the robustness of the symmetry but also for gaining a deeper understanding of quantum black holes. and examine the robustness.
More precisely, we analyze two mini-superspace models at the classical level and show explicitly that Schrödinger symmetry emerges in both cases:
  • (I)
    Spherically-symmetric static mini-superspace with Maxwell field and cosmological constant,
  • (II)
    Spherically-symmetric static mini-superspace with $n$ massless scalar fields.
Model I describes (anti-)de Sitter Reissner-Nordström (AdS-RN) solutions, which reduce to the Schwarzschild solutions in the limit of vanishing electric charge and cosmological constant. It thus serves as a natural first step toward incorporating matter into black hole mini-superspace models. Model II leads to the Janis–Newman–Winicour (JNW) solution Janis:1968zz , which asymptotes to the Schwarzschild geometry at $r\to\infty$ but lacks an event horizon and instead possesses a naked singularity, and it also provides a basic setup for relational formalism using a scalar field as a “clock” Ashtekar:2011ni ; Domagala:2010bm .
In a spherically symmetric spacetime, the areal radius $r$ is well-defined, and it should be natural to consider foliations of spacetime in terms of $r$ or an associated parameter $\tau$. Then, integrating over time (due to the static nature) and over spherical directions (due to spherical symmetry) leaves only $\tau$. This parameter serves as “time”, and we obtain the Hamiltonian $H=N\mathcal{H}$. Here, $N$ is the lapse function, and $\mathcal{H}$ is the Hamiltonian constraint that guarantees the reparametrization invariance for $\tau$. Thus, the two mini-superspace models are a kind of one-dimensional mechanical system with $H=0$ (“black-hole mechanics” Cavaglia:1994yc ; Bodendorfer:2019cyv ; Achour:2021dtj ; BenAchour:2022fif ; BenAchour:2023dgj ; Perez:2023jrq ; Livine:2025soz ). We study the symmetry of such systems.
In Sec.II, we see that for a free-particle system, Schrödinger symmetry can be obtained easily by calculating the algebra of the Noether charges. Here, symmetry means that a Hamiltonian does not change by the transformation (while a more physical meaning is discussed below). Our strategy for obtaining the symmetry in the mini-superspace models is to choose the lapse function $N$ properly and perform canonical transformations so that the Hamiltonian $H$ takes the same form as that of the free particle system. Then, the Schrödinger algebra can be constructed (without imposing $H=0$), and the Schrödinger symmetry appears in terms of the new canonical variables. Here, the solution configuration is represented as a uniform linear motion in the mini-superspace with the momentum constrained by $H=0$, which can be translated into the spacetime geometry through the canonical transformation.
In Sec.III, we implement this procedure explicitly in the spherically-symmetric static vacuum mini-superspace model with cosmological constant and demonstrate the usefulness and consistency by reproducing the 2D-Schrödinger symmetry BenAchour:2023dgj in a systematic manner.
In Sec.IV, we analyze Model I, where the presence of the Maxwell field $A_{0}(r)$ increases the dimension of the mini-superspace by one. By applying our strategy, we show that the 3D Schrödinger symmetry appears. The 2D Schrödinger symmetry is embedded as a subalgebra, and in the limit of vanishing the electric charge, the 3D symmetry reduces to the 2D one. (A)dS-RN metrics can be obtained as the solution in a consistent manner.
In Sec.V, we study Model II, where the dimension of the mini-superspace is $n+2$. By performing a similar analysis but using a different choice of $N$ and a different coordinate in the mini-superspace, we prove that (n+2)(n+2)D Schrödinger symmetry is obtained. If the scalar field contribution is dropped, the 2D symmetry is reproduced. It is remarkable here that the 2D symmetry obtained in the matter decoupling limit of the two models is represented in terms of different lapse functions and mini-superspace coordinates. This is a signal of the covariance of the 2D Schrödinger symmetry in the mini-superspace, which should be an interesting result supporting the robustness of the symmetry. Solving Model II leads to the JNW solution (generalized to $n$ scalar fields) and its “interior” solution (a Kantowski-Sachs-type closed universe) in a unified manner.
In Sec.VI, we discuss the physical meaning of the obtained symmetry under the Hamiltonian constraint $H=0$. First, the generators $G$ of the Schrödinger symmetry obtained for a certain choice of $N$ satisfy $\{G,H\}+\partial_{\tau}G=0$. Momenta $p_{i}$ and angular momenta $j_{ij}$ commute with $H$, and we can show that such generators transform a solution to another one satisfying $H=0$ (dynamical symmetry). As examples, we construct symmetry transformations that map the RN solution to the Schwarzschild one and those that map the JNW solution to the Schwarzschild one. On the other hand, Galilean boost $b^{i}$, dilatation $q_{0}$, and special conformal transformation $q_{-}$ do not commute with $H$, which prevents interpretation as a dynamical symmetry consistent with $H=0$. So, we suggest one possible interpretation: such a non-commutative symmetry transforms the theory itself, $H\to H_{\rm New}$, and the transformed configuration is a physical solution of the new theory with $H_{\rm New}=0$. We will demonstrate this concretely: in Model I, as the effect of adding a two-dimensional cosmological term; in Model II, as the effect of adding a new scalar field. This provides one interpretation of Noether’s theorem in a theory with the Hamiltonian constraint $H\approx 0$ and symmetry generators $G$ that do not commute with $H$.
In Sec.VII, we conclude with future prospects.
Our main result, that Schrodinger symmetries appear in mini-superspace models with different matter contents and different dimensions, should support the robustness of the emergence of Schrodinger symmetry in a fluid limit of (quantum) gravity and contribute to future studies for quantum dynamics of gravity and matter.
II Classical Schrödinger symmetry
II.1 Free particle
We review Schrödinger symmetry Niederer:1972zz ; Duval:2024eod in a classical free particle system and see how to find Schrödinger symmetry in a heuristic manner. We consider
$S[x^{i}(t)]=\int dt\frac{1}{2}g_{ij}\dot{x}^{i}\dot{x}^{j}=\int dtL,$ (1)
where $\dot{x}^{i}\equiv\frac{dx^{i}}{dt}$ and $g_{ij}$ is an invertible symmetric constant metric on the configuration space spanned by $\{x^{i}\}_{i=1}^{d}$, say, $g_{ij}=\delta_{ij}$. Defining the canonical momentum $p_{i}\equiv\frac{\partial L}{\partial\dot{x}^{i}}=g_{ij}\dot{x}^{j}$, the Hamiltonian is given by
$H=\frac{1}{2}g^{ij}p_{i}p_{j},$ (2)
and the canonical relation in the Poisson bracket holds:
$\left\{x^{i},p_{j}\right\}=\delta^{i}_{j},~~\left\{x^{i},x^{j}\right\}=0=\left\{p_{i},p_{j}\right\}$ (3)
Solving the canonical equations
$\dot{x}^{i}=\{x,H\}=g^{ij}p_{j},~~\dot{p}_{i}=\{p_{i},H\}=0,$ (4)
we can obtain the solution, a uniform linear motion in the configuration space:
$x^{i}(t)=g^{ij}\alpha_{j}t+\beta^{i}~~{\rm with}~~p_{i}(t)=\alpha_{i},$ (5)
where $\alpha_{i}$ and $\beta^{i}$ are $2d$ integration constants fixed at $t=0$ that characterize the solution.
Generally, in the canonical form of Noether’s theorem, a symmetry transformation is generated by a function $G(q,p,t)$ that satisfies, in off-shell,
$\{G,H\}+\frac{\partial G}{\partial t}=0,$ (6)
which is the condition for the Hamiltonian to be invariant. [^2] If no Hamiltonian constraint exists, this leads to dynamical symmetries. See Sec.VI for more details. That is, a trajectory generated by such a function $G$ from a solution $(x^{i}(t),p_{i}(t))$,
$x^{\prime i}(t)$ $=x^{i}(t)+\epsilon\{x^{i},G\}|_{x(t),p(t)},$
$p^{\prime}_{i}(t)$ $=p_{i}(t)+\epsilon\{p_{i},G\}|_{x(t),p(t)},$ (7)
also satisfies the canonical equation as another solution with different values of integration constants, where $\epsilon$ is a small parameter. On the other hand, such a generator $G$ is conserved in on-shell:
$\left.\frac{dG}{dt}\right|_{x(t),p(t)}=\left[\left\{G,H\right\}+\frac{\partial G}{\partial t}\right]_{x(t),p(t)}=0.$ (8)
Generally, in the case of the free-particle Hamiltonian, symmetry generators $G$ transform a uniform linear motion with certain values of $(\alpha_{i},\beta^{i})$ to another one with different ones $(\alpha^{\prime}_{i},\beta^{\prime i})$. (We will provide a more explicit argument in Sec.VI.) In particular, the momentum $p_{i}$ and the Galilean boost generator $b^{i}$ defined by
$b^{i}:=x^{i}-g^{ij}p_{j}t$ (9)
satisfy the symmetry condition (6) and play roles of the symmetry generators. Note that using the solution trajectory (5), we have a on-shell relation:
$p_{i}|_{x(t),p(t)}=\alpha_{i},~~b^{i}|_{x(t),p(t)}=\beta^{i},$ (10)
meaning that the initial value of the position, $\beta^{i}$, corresponds to the boost generator $b^{i}$ in on-shell.
Now, quantities consisting of $(b^{i},p_{i})$ also satisfy the symmetry condition (6) through the Leibniz rule. In particular, we can construct the quadratic ones:
$j_{ij}:=g_{ik}b^{k}p_{j}-g_{jk}b^{k}p_{i}=g_{ik}x^{k}p_{j}-g_{jk}x^{k}p_{i},$ (11a)
$q_{+}:=\frac{1}{2}g^{ij}p_{i}p_{j}=H,$ (11b)
$q_{0}:=\frac{1}{2}b^{i}p_{i}=\frac{1}{2}x^{i}p_{i}-Ht,$ (11c)
$q_{-}:=\frac{1}{2}g_{ij}b^{i}b^{j}=\frac{1}{2}g_{ij}x^{i}x^{j}-x^{i}p_{i}t+Ht^{2}.$ (11d)
Noting that $b^{i}$ and $p_{i}$ satisfy
$\left\{b^{i},p_{j}\right\}=\delta^{i}_{j},$ (12)
we can check the following algebra:
$\{b^{i},b^{j}\}=0=\{p_{i},p_{j}\},$ (13a)
$\left\{j_{ij},b^{k}\right\}=\delta^{k}_{i}g_{jl}b^{l}-\delta^{k}_{j}g_{il}b^{l},$ (13b)
$\left\{j_{ij},p_{k}\right\}=g_{ik}p_{j}-g_{jk}p_{i},$ (13c)
$\left\{j_{ij},j_{kl}\right\}=g_{ik}j_{jl}-g_{jk}j_{il}-g_{il}j_{jk}+g_{jl}j_{ik}.$ (13d)
These together with (12) form the Galilean algebra for $g_{ij}=\delta_{ij}$. On the other hand, we can show from (12) that $q_{+},q_{0}$ and $q_{-}$ satisfy $\mathfrak{sl}(2,\mathbb{R})$ algebra:
$\left\{q_{0},q_{\pm}\right\}=\pm q_{\pm},$ (14a)
$\left\{q_{+},q_{-}\right\}=-2q_{0}.$ (14b)
Specifically, $q_{+}$ generates time translations, $q_{0}$ dilatations (scaling transformations), and $q_{-}$ special conformal transformations. These two algebras are intertwined as
$\left\{q_{0},b^{i}\right\}=-\frac{1}{2}b^{i},\quad\left\{q_{0},p_{i}\right\}=\frac{1}{2}p_{i},$ (15a)
$\left\{q_{+},b^{i}\right\}=-g^{ij}p_{j},\quad\left\{q_{-},p_{i}\right\}=g_{ij}b^{j}.$ (15b)
Thus, the full algebra consisting of (12), (13), (14), and (15) forms the $d$-dimensional Schrödinger algebra:
$\mathfrak{sh}(d)=\left(\mathfrak{sl}(2,\mathbb{R})\oplus\mathfrak{so}(d)\right)\oplus_{s}\left(\mathbb{R}^{d}\oplus\mathbb{R}^{d}\right),$ (16)
for $g_{ij}=\delta_{ij}$, where $\oplus_{s}$ denotes a semidirect sum defined by the relations (13b),(13c) and (15). The corresponding symmetry is the d-dimensional Schrödinger symmetry:
$\mathrm{Sh}(d)=\left(\mathrm{SL}(2,\mathbb{R})\times\mathrm{SO}(d)\right)\ltimes\left(\mathbb{R}^{d}\times\mathbb{R}^{d}\right).$ (17)
When the configuration space has a Lorentzian signature, for example,
$g_{ij}=\begin{pmatrix}\begin{matrix}0&1\\ 1&0\end{matrix}&0\\ 0&I_{d-2}\end{pmatrix}\quad\text{or}\quad\begin{pmatrix}\begin{matrix}-1&0\\ 0&1\end{matrix}&0\\ 0&I_{d-2}\end{pmatrix}$, (18)
the generators $j_{ij}$ form the Lie algebra $\mathfrak{so}(1,d-1)$, and the Schrödinger algebra becomes
$\mathfrak{sh}(d)=\left(\mathfrak{sl}(2,\mathbb{R})\oplus\mathfrak{so}(1,d-1)\right)\oplus_{s}\left(\mathbb{R}^{d}\oplus\mathbb{R}^{d}\right),$ (19)
corresponding to the symmetry group:
$\mathrm{Sh}(d)=\left(\mathrm{SL}(2,\mathbb{R})\times\mathrm{SO}(1,d-1)\right)\ltimes\left(\mathbb{R}^{d}\times\mathbb{R}^{d}\right).$ (20)
Generically, gravitational (mini-)superspaces have a Lorentizan signature DeWitt:1967yk . Indeed, a Lorentzian version of the Schrödinger algebra will appear in our models below.
Before proceeding to the next subsection, we emphasize two important points. The first one is that, in the classical theory, not all generators of the Schrödinger symmetry are independent BenAchour:2023dgj . The classical phase space has $2d$ degrees of freedom corresponding to the integration constants $(\alpha_{i},\beta^{i})$, while the number of the generators $b^{i},p_{i},j_{ij},q_{0},q_{\pm}$ is $2d+\frac{1}{2}d(d-1)+3$, with the Casimirs of the symmetry algebra vanishing. In quantum theory, on the other hand, the “shape” of wave functions, corresponding to an infinite number of degrees of freedom, has the physical information of the system and leads to non-vanishing Casimirs of the algebra of the Schrödinger symmetry, which gives the maximal symmetry of the free Schrödinger equation Niederer:1972zz . Therefore, this symmetry, which is classically redundant, would provide, once quantized, the necessary and sufficient symmetry to characterize the system.
The second point is that, the (classical) Schrödinger algebra can be constructed solely from quantities like $(b^{i},p_{i})$ that satisfy both the symmetry condition (6) and the canonical relation (12). When such ones are obtained and the Hamiltonian takes the same form as the free one (2), the remaining generators of the Schrödinger algebra can be obtained in the same way as in (11). This observation provides a powerful guiding principle for finding the Schrödinger algebra in a more generic setup, which will be explored below.
II.2 Canonical-transformation Method
An interesting and important question for understanding the robustness of the emergent symmetry of gravitational systems is: under which condition can the Schrödinger symmetry appear in general? Although finding a complete answer would be non-trivial, there is one related result based on Eisenhart–Duval (ED) lift method; one considers a “lifted” configuration space endowed with a lifted metric and can show that when the lifted space is conformally flat, the Schrödinger symmetry is obtained cccc1220-d9a2-3162-a3c5-a8782bbc8c8b ; universe9120503. Instead of using this method, we will consider more general (but not completely general yet) cases in which the ED method may not be useful, and find a sufficient condition for the Schrödinger symmetry to appear.
Our idea follows from the observation discussed at the end of the previous subsection. If the Hamiltonian takes the same form as that of a free particle, the Schrödinger symmetry is naturally obtained. Therefore, even for systems with non-trivial metrics $g_{ij}(x)$ or potentials $V(x)$, if a canonical transformation exists that transforms the Hamiltonian into the form of the free particle Hamiltonian, the Schrödinger symmetry is automatically derived for the new canonical variables by applying the method in the previous subsection.
More explicitly, we consider a system with a $x$-dependent metric $g_{ij}(x)$ and a potential $V(x)$:
$H=\frac{1}{2}g^{ij}(x)p_{i}p_{j}+V(x),$ (21)
Asuume that we can find a canonical transformation $(x^{i},p_{i})\to(\tilde{x}^{i},\tilde{p}_{i})$ such that this Hamiltonian becomes
$H=\frac{1}{2}\tilde{g}^{ij}\tilde{p}_{i}\tilde{p}_{j},$ (22)
where $\tilde{g}^{ij}$ is a constant metric. Note here that canonical transformations preserve the symplectic structure: $\{x^{i},p_{j}\}=\delta^{i}_{j}\to\{\tilde{x}^{i},\tilde{p}_{j}\}=\delta^{i}_{j}$. Therefore, we can apply the same procedure as in the previous subsection to construct the generators of the Schrödinger algebra. Formally, they are given by replacing $(x^{i},p_{i})$ with $(\tilde{x}^{i},\tilde{p}_{i})$ in the generators (9) and (11):
$\tilde{b}^{i}:=\tilde{x}^{i}-\tilde{g}^{ij}\tilde{p}_{j}t,$ (23a)
$\tilde{j}_{ij}:=\tilde{g}_{ik}\tilde{x}^{k}\tilde{p}_{j}-\tilde{g}_{jk}\tilde{x}^{k}\tilde{p}_{i}=\tilde{g}_{ik}\tilde{b}^{k}\tilde{p}_{j}-\tilde{g}_{jk}\tilde{b}^{k}\tilde{p}_{i},$ (23b)
$\tilde{q}_{+}:=\frac{1}{2}\tilde{g}^{ij}\tilde{p}_{i}\tilde{p}_{j}=H,$ (23c)
$\tilde{q}_{0}:=\frac{1}{2}\tilde{b}^{i}\tilde{p}_{i},$ (23d)
$\tilde{q}_{-}:=\frac{1}{2}\tilde{g}_{ij}\tilde{b}^{i}\tilde{b}^{j}.$ (23e)
By construction, these charges satisfy the symmetry condition (6) and obey the same algebra as in (12), (13), (14), and (15). The above is our strategy for identifying Schrödinger symmetries in several non-trivial models.
[…]
$p_{i}=\frac{\partial F}{\partial x^{i}}.$ (24)
Then, we gain $\tilde{x}^{i}$ by
$\tilde{x}^{i}=\frac{\partial F}{\partial\tilde{p}_{i}}.$ (25)
[…]
Before moving on to applying this method to models, we discuss its subtle points. This is a heuristic approach; whether the desired canonical transformation exists can only be checked by actually performing the analysis. In particular, the first step above is very non-trivial. So, under what general conditions does such a canonical transformation exist? What does its existence physically imply? Understanding these questions would require further studies. Another aspect is that the existence of such a canonical transformation is merely one sufficient condition for the existence of Schrödinger symmetries. How much broader is its applicability compared to the ED-lift method? What is the relationship between these two? We leave these as future challenges.
III Vacuum with Cosmological Constant
To demonstrate the usefulness and consistency of our method, we consider a case where the ED method cannot be applied directly to obtain Schrödinger symmetries, and use our strategy to reproduce a known result. That is, we apply the canonical-transformation method to the spherically-symmetric static vacuum mini-superspace model with cosmological constant $\Lambda$ and rederive the 2D-Schrödinger symmetry BenAchour:2023dgj .
III.1 X-coordinate in the mini-superspace
We first construct the Hamiltonian. We consider a radial foliation of a 4D spherically-symmetric static spacetime, $\mathcal{M}=\mathbb{R}\times\Sigma$, and set the metric as
$ds^{2}=\epsilon\left(-N(\tau)^{2}d\tau^{2}+\frac{4X^{+}(\tau)}{X^{-}(\tau)}dt^{2}\right)+\frac{l_{s}^{2}}{4}{X^{-}(\tau)}^{2}d\Omega^{2}.$ (26)
Here, $\mathbb{R}$ is a domain of the radial variable $\tau$; $\Sigma$ is a 3D hypersurface specified by fixing $\tau$, where $\Sigma$ is timelike for $\epsilon=-1$ and spacelike for $\epsilon=1$; $N(\tau)$ is the lapse function associated with the reparametrization invariance of the radial coordinate; and $l_{s}$ is a fiducial length scale that introduces the dimensionful unit of the 2-sphere radius.
By using this metric, we evaluate the Einstein-Hilbert action with cosmological constant $\Lambda$:
$S_{g}[g_{\mu\nu}]=\frac{1}{16\pi G\hbar}\int_{\mathcal{M}}d^{4}x\sqrt{-g}(R-2\Lambda),$ (27)
Here, for convenience, we introduce the Planck constant $\hbar$ to make dimensional analysis easier. To make the hypersurface $\Sigma$ finite, we set an IR cutoff scale for the range of $t$ integral: $l_{0}\equiv\int^{t_{1}}_{t_{0}}dt=t_{1}-t_{0}$. We then integrate this action over time and 2-sphere, perform integration by parts, and drop boundary terms to obtain the reduced Lagrangian $L_{g}$ via $S_{g}=\int d\tau L_{g}$:
$L_{g}$ $=N\sqrt{\frac{X^{+}}{X^{-}}}\left(-\epsilon cl_{p}\frac{X^{-}}{4X^{+}}\frac{{\dot{X}^{+}}{\dot{X}^{-}}}{N^{2}}+\frac{cl_{p}}{l_{s}^{2}}-\frac{cl_{p}}{l_{\Lambda}^{2}}(X^{-})^{2}\right).$ (28)
Here, $\dot{X}^{+}\equiv\frac{dX^{+}}{d\tau}$, $l_{\Lambda}\equiv\frac{2}{\sqrt{\Lambda}}$, and
$c:=\frac{l_{0}l_{s}^{2}}{l_{p}^{3}}$ (29)
is the dimensionless constant, the ratio between the UV cutoff scale $l_{p}\equiv\sqrt{\hbar G}$ and the IR cutoff scale $l_{0}$.
We introduce the canonical momenta as
$P_{\pm}:=\frac{\partial L_{g}}{\partial{\dot{X}^{\pm}}}=-\frac{\epsilon cl_{p}}{4}\sqrt{\frac{X^{-}}{X^{+}}}\frac{{\dot{X}^{\mp}}}{N},$ (30)
so that the canonical relation holds:
$\{X^{a},P_{b}\}=\delta^{a}_{b},~~~\{X^{a},X^{b}\}=0=\{P_{a},P_{b}\},$ (31)
where $a,b=+,$-, and the Poisson bracket is defined as
$\{f_{1},f_{2}\}\equiv\frac{\partial f_{1}}{\partial X^{a}}\frac{\partial f_{2}}{\partial P_{a}}-\frac{\partial f_{2}}{\partial X^{a}}\frac{\partial f_{1}}{\partial P_{a}},$ (32)
for functions on the phase space: $f_{1}(X,P),~f_{2}(X,P)$. We then obtain the Hamiltonian:
$H_{g}:=$ $P_{+}\dot{X}^{+}+P_{-}\dot{X}^{-}-L_{g}$
$=$ $N\sqrt{\frac{X^{+}}{X^{-}}}\left(-\frac{4}{\epsilon cl_{p}}P_{+}P_{-}-\frac{cl_{p}}{l_{s}^{2}}+\frac{cl_{p}}{l_{\Lambda}^{2}}(X^{-})^{2}\right)$
$\equiv$ $N\mathcal{H}_{g}.$ (33)
$\mathcal{H}_{g}$ is the Hamiltonian constraint, which must be set to zero in order to gain the physical solution: ${\cal H}_{g}\approx 0$. Here, $\approx$ is the weak equality that holds after calculating Poisson brackets, as usual.
We now fix the gauge as
$N(\tau)=\sqrt{\frac{X^{-}}{4X^{+}}}.$ (34)
Then, the Hamiltonian (33) becomes
$H_{g}=$ $-\frac{2}{\epsilon cl_{p}}P_{+}P_{-}-\frac{cl_{p}}{2l_{s}^{2}}+\frac{cl_{p}}{2l_{\Lambda}^{2}}(X^{-})^{2}.$ (35)
This can be expressed as the form of (21):
$H_{g}=\frac{1}{2}\mathcal{G}^{ab}P_{a}P_{b}+\mathcal{V}_{\Lambda}(X)$ (36)
with the quadratic potential $\mathcal{V}_{\Lambda}(X)=-\frac{cl_{p}}{2l_{s}^{2}}+\frac{cl_{p}}{2l_{\Lambda}^{2}}(X^{-})^{2}$ and the metric of the 2D mini-superspace (supermetric)
$\mathcal{G}_{ab}=\begin{pmatrix}0&-\frac{\epsilon cl_{p}}{2}\\ -\frac{\epsilon cl_{p}}{2}&0\end{pmatrix}\quad\text{and}\quad\mathcal{G}^{ab}=\begin{pmatrix}0&-\frac{2}{\epsilon cl_{p}}\\ -\frac{2}{\epsilon cl_{p}}&0\end{pmatrix}$, (37)
where $a,b=+,$-, and the mini-superspace is “Lorentzian”, $\rm{det}\mathcal{G}_{ab}<0$, as consistent with the general result DeWitt:1967yk . Note here that the lifted space is not conformally flat due to $\Lambda\neq 0$, which does not allow one to apply the ED method and directly find Schrodinger symmetries universe9120503.
III.2 2D Schrödinger Symmetry
[…]
According to the explanation in the previous subsection, our first step is to find conserved quantities $f_{a}(X,P)$ and identify them with new momenta $\widetilde{P}_{a}(X,P)$. Noting that $\{H_{g},H_{g}\}=\{H_{g},P_{+}\}=0$, we can write (35) as
$H_{g}$ $=-\frac{2}{\epsilon cl_{p}}P_{+}\left[P_{-}-\frac{\epsilon c^{2}l_{p}^{2}}{4l_{\Lambda}^{2}}\frac{(X^{-})^{2}}{P_{+}}\right]-\frac{cl_{p}}{2l_{s}^{2}}$ (38)
and notice that the bracket part $\cdots$] commutes with $H_{g}$. Therefore, we set
$\widetilde{P}_{+}=P_{+},~~~\widetilde{P}_{-}=P_{-}-\frac{\epsilon c^{2}l_{p}^{2}}{4l_{\Lambda}^{2}}\frac{(X^{-})^{2}}{P_{+}}.$ (39)
[…]
$=X^{+}\widetilde{P}^{+}+X_{-}\widetilde{P}^{-}+\frac{\epsilon c^{2}l_{p}^{2}}{12l_{\Lambda}^{2}}\frac{(X^{-})^{3}}{\widetilde{P}_{+}}$. (40)
Then, the new position coordinates are given through $\widetilde{X}^{a}=\frac{\partial F_{g}}{\partial\widetilde{P}_{a}}$ (25) by
$\widetilde{X}^{+}=X^{+}-\frac{\epsilon c^{2}l_{p}^{2}}{12l_{\Lambda}^{2}}\frac{(X^{-})^{3}}{(P_{+})^{2}},~~~\widetilde{X}^{-}=X^{-}$ (41)
We can check by explicit calculation that these new variables satisfy
$\{\widetilde{X}^{a},\widetilde{P}_{b}\}=\delta^{a}_{b},~~\{\widetilde{X}^{a},\widetilde{X}^{b}\}=0=\{\widetilde{P}_{a},\widetilde{P}_{b}\},$ (42)
as expected from the general property of canonical transformations. Therefore, the Poisson brackets (32) can be calculated in terms of $(\widetilde{X}^{a},\widetilde{P}_{a})$, instead of $(X^{a},P_{a})$.
Thus, we complete the canonical transformation, and the Hamiltonian (35) becomes, in the gauge (34), the desired form:
$H_{g}$ $=-\frac{2}{\epsilon cl_{p}}\widetilde{P}_{+}\widetilde{P}_{-}-\frac{cl_{p}}{2l_{s}^{2}}$ (43)
$=\frac{1}{2}\mathcal{G}^{ab}\widetilde{P}_{a}\widetilde{P}_{b}-\frac{cl_{p}}{2l_{s}^{2}},$
where $\mathcal{G}^{ab}$ is still given by (37).
Now that the Hamiltonian takes the same form as the free-particle Hamiltonian, we can apply the strategy in Sec.II.2 and obtain the Schrödinger symmetry. The set of the generators consists of $\widetilde{P}_{\pm}$ and
$\widetilde{B}^{\pm}:=\widetilde{X}^{\pm}-\mathcal{G}^{\pm a}\widetilde{P}_{a}\tau,$ (44a)
[…]
$\widetilde{Q}_{+}:=\frac{1}{2}\mathcal{G}^{ab}\widetilde{P}_{a}\widetilde{P}_{b}=H_{g}+\frac{cl_{p}}{2l_{s}^{2}},$ (44c)
$\widetilde{Q}_{0}:=\frac{1}{2}\widetilde{B}^{a}\widetilde{P}_{a},$ (44d)
$\widetilde{Q}_{-}:=\frac{1}{2}\mathcal{G}_{ab}\widetilde{B}^{a}\widetilde{B}^{b},$ (44e)
which satisfy the symmetry condition (6) for the Hamiltonian (43). These have the relations (12), (13), (14), and (15), establishing the 2D Schrödinger algebra:
$\mathfrak{sh}(2)=\left(\mathfrak{sl}(2,\mathbb{R})\oplus\mathfrak{so}(1,1)\right)\oplus_{s}\left(\mathbb{R}^{2}\oplus\mathbb{R}^{2}\right).$ (45)
Its exponentiation gives the 2D Schrödinger group,
$\mathrm{Sh}(2)=\left(\mathrm{SL}(2,\mathbb{R})\times\mathrm{SO}(1,1)\right)\ltimes\left(\mathbb{R}^{2}\times\mathbb{R}^{2}\right).$ (46)
Thus, we have applied the canonical-transformation method and reproduced systematically the 2D Schrödinger symmetry, the result obtained in a heuristic manner BenAchour:2023dgj .
We conclude this derivation with three remarks. First, the above symmetry holds for the specific choice of the the lapse function (34). Second, the Hamiltonian constraint is not used to derive the symmetry. Third, the constraint itself appears as one of the generators, $\widetilde{Q}_{+}=H_{g}+\frac{cl_{p}}{2l_{s}^{2}}$. These three are intertwined in understanding the physical meaning of the symmetry (see Sec.VI).
III.3 (Anti)de-Sitter Schwarzschild metric
Before closing this section, we here show that the solution of the Hamiltonian (35) is uniquely given by the (anti)de Sitter Schwarzschild metric, which provides a physical aspect of this system. In the derivation, we impose the Hamiltonian constraint $H_{g}\approx 0$ as late as possible to make its role clear.
First, the canonical equations provide a uniform linear motion in the configuration space $\{\widetilde{X}^{a}\}$:
$\dot{\widetilde{P}_{\pm}}$ $=\{\widetilde{P}_{\pm},H_{g}\}=0\Rightarrow\widetilde{P}_{\pm}(\tau)=\rm{const.}$
$\dot{\widetilde{X}^{\pm}}$ $=\{\widetilde{X}^{\pm},H_{g}\}=-\frac{2}{\epsilon cl_{p}}\widetilde{P}_{\mp}\Rightarrow\widetilde{X}^{\pm}(\tau)=-\frac{2}{\epsilon cl_{p}}\widetilde{P}_{\mp}\tau+\widetilde{B}^{\pm}$. (47)
Here and in the remainder of this section, we regard $\widetilde{B}^{\pm}$ and $\widetilde{P}_{\pm}$ as constants $\beta^{\pm}$ and $\alpha_{\pm}$, respectively.
We note from the metric (26) that the areal radius of the 2-sphere, $r$, should be identified as
$r=\frac{l_{s}}{2}X^{-}.$ (48)
Therefore, from the canonical transformation (41) and the trajectory (III.3), we can have $\frac{2}{l_{s}}r=X^{-}(\tau)=\widetilde{X}^{-}(\tau)=-\frac{2}{\epsilon cl_{p}}\widetilde{P}_{+}\tau+\widetilde{B}^{-}$, leading to
$\tau=-\frac{\epsilon\bar{c}}{\widetilde{P}_{+}}\left(r-\frac{l_{s}}{2}\widetilde{B}^{-}\right)\Rightarrow d\tau=-\frac{\epsilon\bar{c}}{\widetilde{P}_{+}}dr,$ (49)
where we define
$\bar{c}\equiv\frac{cl_{p}}{l_{s}}=\frac{l_{0}l_{s}}{l_{p}^{2}}.$ (50)
Similarly, we have for $X^{+}$
$X^{+}=\frac{\widetilde{P}_{-}}{\widetilde{P}_{+}}X^{-}+\frac{1}{\widetilde{P}_{+}}(\widetilde{B}^{+}\widetilde{P}_{+}-\widetilde{B}^{-}\widetilde{P}_{-})+\frac{\epsilon c^{2}l_{p}^{2}}{12l_{\Lambda}^{2}}\frac{(X^{-})^{3}}{\widetilde{P}_{+}^{2}}$ (51)
Combining these and the gauge condition (34), we can calculate
$-\epsilon N^{2}d\tau^{2}$ $=-\epsilon\frac{X^{-}}{4X^{+}}\frac{\bar{c}^{2}}{\widetilde{P}_{+}^{2}}dr^{2}$
$=\frac{1}{\frac{4\widetilde{P}_{+}\widetilde{P}_{-}}{-\epsilon\bar{c}^{2}}-\frac{2GM}{r}-\frac{\Lambda}{3}r^{2}}dr^{2}$
$\approx\frac{1}{1-\frac{2GM}{r}-\frac{\Lambda}{3}r^{2}}dr^{2},$ (52)
where we have defined
$M\equiv\frac{\epsilon l_{s}}{G\bar{c}^{2}}\widetilde{P}_{+}(\widetilde{P}_{+}\widetilde{B}^{+}-\widetilde{P}_{-}\widetilde{B}^{-})=\frac{2}{G\bar{c}^{3}}\widetilde{P}_{+}\widetilde{J}_{+-}$, (53)
and applied the Hamiltonian constraint condition in the last step: (from (43))
$H_{g}\approx 0\Rightarrow\widetilde{P}_{+}\widetilde{P}_{-}\approx-\frac{\epsilon\bar{c}^{2}}{4}.$ (54)
Similarly, we can have
$\epsilon\frac{4X^{+}}{X^{-}}dt^{2}$ $=\epsilon\frac{4X^{+}}{X^{-}}\frac{\widetilde{P}_{+}^{2}}{\bar{c}^{2}}\frac{\bar{c}^{2}}{\widetilde{P}_{+}^{2}}dt^{2}$
$=-\left(\frac{4\widetilde{P}_{+}\widetilde{P}_{-}}{-\epsilon\bar{c}^{2}}-\frac{2GM}{r}-\frac{\Lambda}{3}r^{2}\right)dt_{\infty}^{2}$
$\approx-\left(1-\frac{2GM}{r}-\frac{\Lambda}{3}r^{2}\right)dt_{\infty}^{2},$ (55)
where we have defined the time at $r=\infty$ as $dt_{\infty}\equiv\frac{\bar{c}}{\widetilde{P}_{+}}dt$. Thus, after solving the canonical equations and imposing $H_{g}\approx 0$, the metric (26) becomes the (anti)de-Sitter Schwarzschild metric:
$ds^{2}\approx$ $-\left(1-\frac{2GM}{r}-\frac{\Lambda}{3}r^{2}\right)dt_{\infty}^{2}$
$+\left(1-\frac{2GM}{r}-\frac{\Lambda}{3}r^{2}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}.$ (56)
Here, the formula (53) plays the role of the energy of the system. Indeed, in the case of $\Lambda=0$ and $\epsilon=-1$, it agrees with the Komar mass Wald:1984rg (see Appendix A). Note that $M$ commutes with $H_{g}$ (43) and thus is gauge invariant.
IV Maxwell Field
In the previous section, we have succeeded in applying the canonical-transformation method to the vacuum mini-superspace with cosmological constant $\Lambda$ and reproducing the known result, the 2D Schrödinger symmetry. In the following sections, we will apply the method to new cases where matter fields exist, increasing the dimension of the mini-superspace. In this section, we will consider Maxwell field $A_{\mu}$ with $\Lambda$ and show that the 3D Schrödinger symmetry appears.
IV.1 3D Schrödinger symmetry
We follow the same steps as in the previous section to derive the symmetry. We first evaluate the Hamiltonian for the reduced Lagrangian of the action
$S_{gE}[g_{\mu\nu},A_{\mu}]$ $=\frac{1}{16\pi G\hbar}\int_{\mathcal{M}}d^{4}x\sqrt{-g}(R-2\Lambda)$
$~~~~~~~~~~-\frac{1}{4\hbar}\int_{\mathcal{M}}d^{4}x\sqrt{-g}F_{\mu\nu}F^{\mu\nu}$ (57)
under the assumption $A_{\mu}(x)dx^{\mu}=A_{0}(\tau)dt$ in the spherically-symmetric metric (26), where $F_{\mu\nu}\equiv\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. Noting $F_{\mu\nu}F^{\mu\nu}=2g^{tt}g^{\tau\tau}\dot{A}_{0}^{2}$ and integrating time and 2-sphere, the Maxwell part $L_{E}$ is given through $S_{E}=\int d\tau L_{E}$ by
$L_{E}=\frac{cl_{p}}{4N}\sqrt{\frac{X^{-}}{X^{+}}}(X^{-})^{2}(\dot{X}^{A})^{2}$ (58)
where we define $X^{A}\equiv\sqrt{\frac{\pi}{\hbar}}l_{p}A_{0}=\sqrt{\pi G}A_{0}$. Introducing the canonical momentum for $X^{A}$ as
$P_{A}:=\frac{\partial L_{E}}{\partial\dot{X}_{A}}=\frac{cl_{p}}{2N}\sqrt{\frac{X^{-}}{X^{+}}}(X^{-})^{2}\dot{X}^{A},$ (59)
we have the Hamiltonian for the Maxwell part:
$H_{E}:=\dot{X}^{A}P_{A}-L_{E}=N\sqrt{\frac{X^{+}}{X^{-}}}\frac{P_{A}^{2}}{cl_{p}(X^{-})^{2}}.$ (60)
Combing this and the gravity one (33), we gain the total Hamiltonian:
$H_{gE}$ $=H_{g}+H_{E}$
$=N\sqrt{\frac{X^{+}}{X^{-}}}\left[-\frac{4}{\epsilon cl_{p}}P_{+}P_{-}+\frac{P_{A}^{2}}{cl_{p}(X^{-})^{2}}\right$.
$~~~~~~~~~~~~~~~~~~~~~~~~\left.+\frac{cl_{p}}{l_{\Lambda}^{2}}(X^{-})^{2}-\frac{cl_{p}}{l_{s}^{2}}\right]$
$\equiv N\mathcal{H}_{gE}.$ (61)
Here, $\mathcal{H}_{gE}$ is the Hamiltonian constraint, which becomes $\mathcal{H}_{gE}\approx 0$ on shell, and the canonical variables $(X^{+},X^{-},X^{A};P_{+},P_{-},P_{A})$ satisfy the relation (31).
In the chosen $N$ (34), the Hamiltonian (IV.1) becomes
$H_{gE}$ $=-\frac{2}{\epsilon cl_{p}}P_{+}P_{-}+\frac{P_{A}^{2}}{2cl_{p}(X^{-})^{2}}+\frac{cl_{p}}{2l_{\Lambda}^{2}}(X^{-})^{2}-\frac{cl_{p}}{2l_{s}^{2}}$ (62)
$=\frac{1}{2}\mathcal{G}^{ab}_{E}P_{a}P_{b}+\mathcal{V}_{\Lambda}(X),$ (63)
where a,b=+,,Aa,b=+,-,A, and the supermetric is 3-dimensional:
$\mathcal{G}_{Eab}$ $=cl_{p}\begin{pmatrix}0&-\frac{\epsilon}{2}&0\\ -\frac{\epsilon}{2}&0&0\\ 0&0&(X^{-})^{2}\end{pmatrix}$,
$\mathcal{G}_{E}^{ab}$ $=\frac{1}{cl_{p}}\begin{pmatrix}0&-\frac{2}{\epsilon}&0\\ -\frac{2}{\epsilon}&0&0\\ 0&0&\frac{1}{(X^{-})^{2}}\end{pmatrix}$, (64)
This is again “Lorentzian”, $\det\mathcal{G}_{Eab}<0$, and the lifted space is not conformally flat.
We next perform a canonical transformation to express the Hamiltonian $H_{gE}$ (62) as the free-particle form. We insert $\frac{1}{2cl_{p}}(-P_{A}^{2}+P_{A}^{2})=0$ into $H_{gE}$ and write it as
$H_{gE}$
$=$ $-\frac{2}{\epsilon cl_{p}}P_{+}\left[P_{-}+\frac{\epsilon P_{A}^{2}}{4P_{+}}\left(1-\frac{1}{(X^{-})^{2}}\right)-\frac{\epsilon c^{2}l_{p}^{2}}{4l_{\Lambda}^{2}}\frac{(X^{-})^{2}}{P_{+}}\right]$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\frac{1}{2cl_{p}}P_{A}^{2}-\frac{cl_{p}}{2l_{s}^{2}}.$ (65)
Noting $\{H_{gE},H_{gE}\}=\{P_{+},H_{gE}\}=\{P_{A},H_{gE}\}=0$, the bracket $\cdots$] also commutes with $H_{gE}$. Therefore, we set the new momenta as
$\widetilde{P}_{+}$ $\equiv P_{+},$
$\widetilde{P}_{-}$ $\equiv P_{-}+\frac{\epsilon P_{A}^{2}}{4P_{+}}\left(1-\frac{1}{(X^{-})^{2}}\right)-\frac{\epsilon c^{2}l_{p}^{2}}{4l_{\Lambda}^{2}}\frac{(X^{-})^{2}}{P_{+}},$
$\widetilde{P}_{A}$ $\equiv P_{A}.$ (66)
[…]
$=X^{+}\widetilde{P}^{+}+X_{-}\widetilde{P}^{-}+X^{A}\widetilde{P}_{A}$
$~~-\frac{\epsilon\widetilde{P}_{A}^{2}}{4\widetilde{P}_{+}}\left(X^{-}+\frac{1}{X^{-}}\right)+\frac{\epsilon c^{2}l_{p}^{2}}{12l_{\Lambda}^{2}}\frac{(X^{-})^{3}}{\widetilde{P}_{+}}$. (67)
In turn, the new coordinates $\widetilde{X}^{a}$ are obtained through $\widetilde{X}^{a}=\frac{\partial F_{gE}}{\partial\widetilde{P}_{a}}$ as
$\widetilde{X}^{+}$ $=X^{+}+\frac{\epsilon P_{A}^{2}}{4P_{+}^{2}}\left(X^{-}+\frac{1}{X^{-}}\right)-\frac{\epsilon c^{2}l_{p}^{2}}{12l_{\Lambda}^{2}}\frac{(X^{-})^{3}}{P_{+}^{2}}$
$\widetilde{X}^{-}$ $=X^{-}$
$\widetilde{X}^{A}$ $=X^{A}-\frac{\epsilon P_{A}}{2P_{+}}\left(X^{-}+\frac{1}{X^{-}}\right).$ (68)
Indeed, these new variables $(\widetilde{X}^{a},\widetilde{P}_{a})$ satisfy the canonical relations (42). We thus reach the desired form:
$H_{gE}$ $=-\frac{2}{\epsilon cl_{p}}\widetilde{P}_{+}\widetilde{P}_{-}+\frac{1}{2cl_{p}}\widetilde{P}_{A}^{2}-\frac{cl_{p}}{2l_{s}^{2}}$ (69)
$=\frac{1}{2}\widetilde{\mathcal{G}}^{ab}_{E}\widetilde{P}_{a}\widetilde{P}_{b}-\frac{cl_{p}}{2l_{s}^{2}}$
with the constant supermetric
$\widetilde{\mathcal{G}}_{Eab}=cl_{p}\begin{pmatrix}0&-\frac{\epsilon}{2}&0\\ -\frac{\epsilon}{2}&0&0\\ 0&0&1\end{pmatrix},~\widetilde{\mathcal{G}}_{E}^{ab}=\frac{1}{cl_{p}}\begin{pmatrix}0&-\frac{2}{\epsilon}&0\\ -\frac{2}{\epsilon}&0&0\\ 0&0&1\end{pmatrix}$. (70)
Finally, we find the symmetry. According to the strategy of Sec.II.2, we set the generators as
$\widetilde{P}_{a},~~\widetilde{B}^{a}:=\widetilde{X}^{a}-\widetilde{\mathcal{G}}^{ab}_{E}\widetilde{P}_{b}\tau,$ (71a)
$\widetilde{J}_{ab}:=\widetilde{\mathcal{G}}_{Eac}\widetilde{X}^{c}\widetilde{P}_{b}-\widetilde{\mathcal{G}}_{Ebc}\widetilde{X}^{c}\widetilde{P}_{a}$
$~~~~~~=\widetilde{\mathcal{G}}_{Eac}\widetilde{B}^{c}\widetilde{P}_{b}-\widetilde{\mathcal{G}}_{Ebc}\widetilde{B}^{c}\widetilde{P}_{a},$ (71b)
$\widetilde{Q}_{+}:=\frac{1}{2}\widetilde{\mathcal{G}}^{ab}_{E}\widetilde{P}_{a}\widetilde{P}_{b}=H_{gE}+\frac{cl_{p}}{2l_{s}^{2}}$ (71c)
$\widetilde{Q}_{0}:=\frac{1}{2}\widetilde{B}^{a}\widetilde{P}_{a},$ (71d)
$\widetilde{Q}_{-}:=\frac{1}{2}\widetilde{\mathcal{G}}_{Eab}\widetilde{B}^{a}\widetilde{B}^{b},$ (71e)
where (a,b,c)=(+,,A)(a,b,c)=(+,-,A). These constitute the 3D-Schrödinger algebra
$\mathfrak{sh}(3)=\left(\mathfrak{sl}(2,\mathbb{R})\oplus\mathfrak{so}(2,1)\right)\oplus_{s}\left(\mathbb{R}^{3}\oplus\mathbb{R}^{3}\right),$ (72)
meaning that the system possesses the 3D-Schrödinger symmetry.
[…]
IV.2 (Anti) de-Sitter Reissner-Nordström metric
We here check that the on-shell metric of this system is uniquely given by the (Anti) de-Sitter Reissner-Nordström metric.
We take the same steps as in Sec.III.3. We first solve the canonical equations (III.3) where $H_{g}$ is replaced with $H_{gE}$ (69). We have the relation (49) between the parameter $\tau$ and the areal radius $r$. As in (51), we can calculate
[…]
$-\frac{\epsilon\widetilde{P}_{A}^{2}}{4\widetilde{P}_{+}^{2}}\left(X^{-}+\frac{1}{X^{-}}\right)+\frac{\epsilon c^{2}l_{p}^{2}}{12l_{\Lambda}^{2}}\frac{(X^{-})^{3}}{\widetilde{P}_{+}^{2}}.$ (73)
We use these ingredients and the lapse (34) and obtain
$-\epsilon N^{2}d\tau^{2}$ $=-\epsilon\frac{X^{-}}{4X^{+}}\frac{\bar{c}^{2}}{\widetilde{P}_{+}^{2}}dr^{2}$
$=\frac{1}{\frac{1}{\bar{c}^{2}}\left(-\frac{4}{\epsilon}\widetilde{P}_{+}\widetilde{P}_{+}+\widetilde{P}_{A}^{2}\right)-\frac{2GM}{r}+\frac{G\mathcal{Q}^{2}}{4\pi r^{2}}-\frac{\Lambda}{3}r^{2}}dr^{2},$
$\approx\frac{1}{1-\frac{2GM}{r}+\frac{G\mathcal{Q}^{2}}{4\pi r^{2}}-\frac{\Lambda}{3}r^{2}}dr^{2}.$ (74)
In the second line, we applied the mass formula (53) and introduced
$\mathcal{Q}\equiv\sqrt{\frac{\pi}{G}}\frac{l_{s}}{\bar{c}}\widetilde{P}_{A},$ (75)
which indeed agrees with the standard definition of electric charge (see Appendix A). In the last line, we employed the Hamiltonian constraint (from (69)):
$H_{gE}\approx 0\Rightarrow-\frac{4}{\epsilon}\widetilde{P}_{+}\widetilde{P}_{-}+\widetilde{P}_{A}^{2}\approx\bar{c}^{2}.$ (76)
Similarly, we have
$\epsilon\frac{4X^{+}}{X^{-}}dt^{2}$ $=\epsilon\frac{4X^{+}}{X^{-}}\frac{\widetilde{P}_{+}^{2}}{\bar{c}^{2}}\frac{\bar{c}^{2}}{\widetilde{P}_{+}^{2}}dt^{2}$
$\approx-\left(1-\frac{2GM}{r}+\frac{G\mathcal{Q}^{2}}{4\pi r^{2}}-\frac{\Lambda}{3}r^{2}\right)dt_{\infty}^{2},$ (77)
where we used $dt_{\infty}\equiv\frac{\bar{c}}{\widetilde{P}_{+}}dt$. Thus, the solution is given uniquely by the AdS Reissner-Nordström metric:
$ds^{2}\approx$ $-\left(1-\frac{2GM}{r}+\frac{G\mathcal{Q}^{2}}{4\pi r^{2}}-\frac{\Lambda}{3}r^{2}\right)dt_{\infty}^{2}$
$+\left(1-\frac{2GM}{r}+\frac{G\mathcal{Q}^{2}}{4\pi r^{2}}-\frac{\Lambda}{3}r^{2}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}.$ (78)
V Multiple Massless Scalar Fields
As another case with matter fields, we discuss the mini-superspace with $n$ massless scalar fields, whose solution corresponds to a generalized Janis–Newman–Winicour (JNW) solution Janis:1968zz and its “interior” Chiou:2008eg , and we show that $(2+n)$ D-Schrödinger symmetry appears. An interesting point here is that the symmetry is obtained by using a different coordinate system $Y^{a}$ in the mini-superspace than the previous one $X^{a}$, together with a different choice of gauge $N$.
V.1 Y-coordinate in the mini-superspace
We consider a coupled system of gravity and $n$ free massless scalar fields $\phi_{I}=\phi_{I}(\tau)$ and construct a Hamiltonian suitable for the canonical-transformation method. Before going to the details, it is interesting to see what would happen if we included mass terms and cosmological constant. So, let us start with a more general action:
$S_{g\phi}[g_{\mu\nu},\phi_{I}]$ $=\frac{1}{16\pi G\hbar}\int d^{4}x\sqrt{-g}(R-2\Lambda)$
$-\frac{1}{2\hbar}\sum_{I=1}^{n}\int d^{4}x\sqrt{-g}\left(g^{\mu\nu}\partial_{\mu}\phi_{I}\partial_{\nu}\phi_{I}+m_{I}^{2}\phi_{I}^{2}\right).$ (79)
First, if we use the ansatz metric (26), which we have considered so far, and evaluate the reduced Hamiltonian $H_{g\phi}$, we can see that it is difficult to apply the canonical-transformation method as in the previous cases. [^4] For the metric (26) and $\phi_{I}=\phi_{I}(\tau)$, the action of scalar fields reduces to $S_{\phi}=\int d\tau L_{\phi}$ with $L_{\phi}=cl_{p}N\sqrt{\frac{X^{+}}{X^{-}}}(X^{-})^{2}\sum_{I}\left(\frac{1}{\epsilon N^{2}}(\dot{X}^{I})^{2}-m_{I}^{2}(X^{I})^{2}\right),$ where $X^{I}\equiv\sqrt{\frac{\pi}{\hbar}}l_{p}\phi_{I}$. Defining the momenta $P_{I}\equiv\frac{\partial L_{\phi}}{\partial\dot{X}^{I}}$, we can calculate the Hamiltonian $H_{\phi}\equiv\sum_{I}P_{I}\dot{X}^{I}-L_{\phi}$ and add the gravity part $H_{g}$ (33) to reach the reduced Hamiltonian $H_{g\phi}=H_{g}+H_{\phi}$: $H_{g\phi}$ $=N\sqrt{\frac{X^{+}}{X^{-}}}\left[-\frac{4}{\epsilon cl_{p}}P_{+}P_{-}-\frac{cl_{p}}{l_{s}^{2}}+\frac{cl_{p}}{l_{\Lambda}^{2}}(X^{-})^{2}\right$. $+\left.\sum_{I}\left(\frac{\epsilon}{4cl_{p}X^{+}X^{-}}P_{I}^{2}+cl_{p}m_{I}^{2}(X^{-})^{2}(X^{I})^{2}\right)\right].$ Here, the mass terms lead to a non-trivial interaction between $X^{-}$ and $X^{I}$. Even if we choose the lapse (34), $N=\sqrt{\frac{X^{-}}{4X^{+}}}$, and set $m_{I}=0$ and $\Lambda=0$, the term of $P_{I}^{2}$ would remain mixing $X^{+}$ and $X^{-}$, and it would be difficult to find two quantities that commute with $H_{g\phi}$.
One idea for overcoming this situation is to employ another coordinate system in the mini-superspace. We consider the $Y$-coordinate system related to the $X$-coordinate system as
$Y^{0}:=\frac{1}{4}\log X^{+}X^{-},~Y^{1}:=\frac{1}{4}\log\frac{X^{+}}{X^{-}},~Y^{I}:=\sqrt{\frac{\pi}{\hbar}}l_{p}\phi_{I},$ (80)
in which the spacetime metric (26) is expressed as
$ds^{2}=\epsilon\left(-N^{2}d\tau^{2}+4e^{4Y^{1}}dt^{2}\right)+\frac{l_{s}^{2}}{4}e^{4Y^{0}-4Y^{1}}d\Omega^{2}.$ (81)
We can use this and evaluate the action (V.1) by performing integration by parts and dropping boundary terms, to obtain the reduced Lagrangian $L_{g\phi}$ via $S_{g\phi}=\int d\tau L_{g\phi}$:
$L_{g\phi}$ $=Ne^{4Y^{0}-2Y^{1}}cl_{p}\left[\frac{1}{\epsilon N^{2}}\left(-(\dot{Y}^{0})^{2}+(\dot{Y}^{1})^{2}+\sum_{I}(\dot{Y}^{I})^{2}\right)\right$.
$~~~~~~~~~~~~~\left.-\sum_{I}m_{I}^{2}(Y^{I})^{2}-\frac{1}{l_{\Lambda}^{2}}+\frac{1}{l_{s}^{2}}e^{-4Y^{0}+4Y^{1}}\right].$ (82)
Defining the momenta $\Pi_{a}\equiv\frac{\partial L_{g\phi}}{\partial\dot{Y}^{a}}$ (for $a=0,1,I$), we have the reduced Hamiltonian:
$H_{g\phi}$ $=Ne^{-4Y^{0}+2Y^{1}}\left[\frac{1}{4\epsilon cl_{p}}\left(-\Pi_{0}^{2}+\Pi_{1}^{2}+\sum_{I}\Pi_{I}^{2}\right)\right$.
$\left.-\frac{cl_{p}}{l_{s}^{2}}e^{4Y^{0}}+cl_{p}e^{8Y^{0}-4Y^{1}}\left(\sum_{I}m_{I}^{2}(Y^{I})^{2}+\frac{1}{l_{\Lambda}^{2}}\right)\right].$ (83)
The mass terms connect all the coordinates $Y^{0},Y^{1},Y^{I}$, and the cosmological constant links $Y^{0}$ and $Y^{1}$, which still makes the application of our method difficult.
In the following, we assume
$m_{I}=0,~~~\Lambda=0$ (84)
and set the lapse as
$N=2e^{4Y^{0}-2Y^{1}}.$ (85)
Then, we reach a simple form of the Hamiltonian:
$H_{g\phi}$ $=\frac{1}{2\epsilon cl_{p}}\left(-\Pi_{0}^{2}+\Pi_{1}^{2}+\sum_{I}\Pi_{I}^{2}\right)-\frac{2cl_{p}}{l_{s}^{2}}e^{4Y^{0}}$ (86)
$=\frac{1}{2\epsilon cl_{p}}\eta^{ab}\Pi_{a}\Pi_{b}-\frac{2cl_{p}}{l_{s}^{2}}e^{4Y^{0}},$
where (diagonal of $\eta^{ab})\equiv(-1,1,\delta_{IJ})$ is a $(2+n)\times(2+n)$ “Lorentzian” constant matrix. This contains only one coordinate $Y^{0}$ so that we can implement the canonical-transformation program easily.
V.2 $(2+n)$ D-Schrödinger symmetry
We now follow our strategy and show the appearance of the $(2+n)$ D-Schrödinger symmetry.
We first change the canonical variables. In the Hamiltonian $H_{g\phi}$ (86), $\Pi_{1}$ and $\Pi_{I}$ commute with $H_{g\phi}$, and therefore, the sum of the remaining two terms must commute with $H_{g\phi}$. This suggests as the new momenta
$\widetilde{\Pi}_{0}\equiv\sqrt{\Pi_{0}^{2}+4\epsilon\bar{c}^{2}e^{4Y^{0}}},~\widetilde{\Pi}_{1}\equiv\Pi_{1},~\widetilde{\Pi}_{I}\equiv\Pi_{I}.$ (87)
A generating function corresponding to these can be identified with
[…]
$=\int dY^{0}\sqrt{\widetilde{\Pi}_{0}^{2}-4\epsilon\bar{c}^{2}e^{4Y^{0}}}+\int dY^{1}\widetilde{\Pi}_{1}+\sum_{I}\int dY^{I}\widetilde{\Pi}_{I}$
$=\frac{1}{2}\sqrt{\widetilde{\Pi}_{0}^{2}-4\epsilon\bar{c}^{2}e^{4Y^{0}}}+\frac{\widetilde{\Pi}_{0}}{4}\log\left(\frac{\epsilon\widetilde{\Pi}_{0}-\epsilon\sqrt{\widetilde{\Pi}_{0}^{2}-4\epsilon\bar{c}^{2}e^{4Y^{0}}}}{\widetilde{\Pi}_{0}+\sqrt{\widetilde{\Pi}_{0}^{2}-4\epsilon\bar{c}^{2}e^{4Y^{0}}}}\right)$
$~~~~~~~~+Y^{1}\widetilde{\Pi}^{1}+\sum_{I}Y_{I}\widetilde{\Pi}^{I}.$ (88)
In turn, this generates the new coordinates,
$\widetilde{Y}^{0}$ $=\frac{1}{4}\log\left(\frac{\epsilon\Pi_{0}-\epsilon\sqrt{\Pi_{0}^{2}-4\epsilon\bar{c}^{2}e^{4Y^{0}}}}{\Pi_{0}+\sqrt{\Pi_{0}^{2}-4\epsilon\bar{c}^{2}e^{4Y^{0}}}}\right),$
$~~~~~~~~~\widetilde{Y}^{1}=Y^{1},~~\widetilde{Y}^{I}=Y^{I},$ (89)
and the canonical relation holds indeed:
$\{\widetilde{Y}^{a},\widetilde{\Pi}_{b}\}=\delta^{a}_{b},~\{\widetilde{Y}^{a},\widetilde{Y}^{b}\}=0=\{\widetilde{\Pi}_{a},\widetilde{\Pi}_{b}\}.$ (90)
Thus, the Hamiltonian (86) is expressed in terms of $(\widetilde{Y}^{a},\widetilde{\Pi}_{a})$ as
$H_{g\phi}$ $=\frac{1}{2(-\epsilon)cl_{p}}\left(-\widetilde{\Pi}_{0}^{2}+\widetilde{\Pi}_{1}^{2}+\sum_{I}\widetilde{\Pi}_{I}^{2}\right)$
$=\frac{1}{2}\mathcal{G}^{ab}_{\phi}\widetilde{\Pi}_{a}\widetilde{\Pi}_{b},$ (91)
with the (2+n)-dimensional constant supermetric:
$\mathcal{G}_{\phi ab}=(-\epsilon)cl_{p}\eta_{ab},~~\mathcal{G}_{\phi}^{ab}=\frac{1}{(-\epsilon)cl_{p}}\eta^{ab},$ (92)
where $\eta_{ab}=\mathrm{diag}(-1,1,\cdots.,1)$. Here, for convenience, we multiplied 1-1 in the right hand side so that the Hamiltonian (V.2) for $\epsilon=-1$ corresponds to the Hamiltonian constraint of a relativistic massless particle, $\mathcal{H}_{\rm{rel}}=-E^{2}+P_{i}^{2}$, where $\widetilde{\Pi}_{0}$ plays a role of “energy EE”, although this is not important for the dynamics since it is determined essentially by $H_{g\phi}\approx 0$ as discussed in Sec.I.
Now that the Hamiltonian (V.2) has exactly same form as the free-particle Hamiltonian (2), we can replace $(x^{i},p_{i},g^{ij})$ with $(\widetilde{Y}^{a},\widetilde{\Pi}_{a},\mathcal{G}^{ab}_{\phi})$ and apply the argument in Sec.II.1 to conclude that the spherically-symmetric static system coupled with $n$ free massless scalar fields has the $(2+n)$ D-Schrödinger symmetry:
$\mathrm{Sh}(2+n)=\left(\mathrm{SL}(2,\mathbb{R})\times\mathrm{SO}(1,1+n)\right)\ltimes\left(\mathbb{R}^{2+n}\times\mathbb{R}^{2+n}\right).$ (93)
The $(2+n)$ D-Schrödinger algebra $\mathfrak{sh}(2+n)$ consists of three sectors. The gravity sector has only gravity components $(\widetilde{\Pi}_{0},\widetilde{\Pi}_{1},\widetilde{B}^{0},\widetilde{B}^{1},\widetilde{J}_{01})$, forming together with $\widetilde{Q}_{\pm},\widetilde{Q}_{0}$ a sub-algebra $\mathfrak{sh}(2)$; the scalar-field sector is composed by $(\widetilde{\Pi}_{I},\widetilde{B}^{I},\widetilde{J}_{IJ})$, leading together with $\widetilde{Q}_{\pm},\widetilde{Q}_{0}$ to a sub-algebra $\mathfrak{sh}(n)$; and the gravity-scalar sector contains the mixed components $\widetilde{J}_{0I},\widetilde{J}_{1I}$. Therefore, in the matter decoupling limit where $\widetilde{\Pi}_{I}\to 0$, only the gravity sector remains, and the algebra reduces to the 2D-Schrödinger algebra $\mathfrak{sh}(2)$.
The remarkable point here is that this algebra $\mathfrak{sh}(2)$ is expressed in terms of $Y$-coordinates with the lapse choice (85) while the one obtained in Sec.III.2 is written in terms of $X$-coordinates under a different gauge (34). Therefore, the 2D-Schrödinger symmetry can be realized in different choices of mini-superspace coordinates and gauges. This means the covariance of the Schrödinger symmetry in the mini-superspace, which signals a robustness of the symmetry (at least, for pure-gravity cases).
V.3 Generalized JNW metric and its “interior”
To see the physics of this system, we solve the Hamiltonian and derive the generalized Janis-Newman-Winicour (JNW) metric with $n$ massless scalar fields and its “interior” metric.
We first solve the canonical equations for $H_{g\phi}$ (V.2):
$\dot{\widetilde{\Pi}}_{a}=\{\widetilde{\Pi}_{a},H_{g\phi}\}=0$ $\Rightarrow\widetilde{\Pi}_{a}(\tau)={\rm const}.$
$\dot{\widetilde{Y}}^{a}=\{\widetilde{Y}^{a},H_{g\phi}\}=\frac{\eta^{ab}\widetilde{\Pi}_{b}}{(-\epsilon)cl_{p}}$ $\Rightarrow\widetilde{Y}^{a}(\tau)=\frac{\eta^{ab}\widetilde{\Pi}_{b}}{(-\epsilon)cl_{p}}\tau+\widetilde{B}^{a},$ (94)
where $(\widetilde{\Pi}_{a},\widetilde{B}^{a})$ are considered as constants characterizing the trajectory again. More explicitly, we have
$\widetilde{Y}^{0}(\tau)=\frac{\widetilde{\Pi}_{0}}{\epsilon cl_{p}}\tau+\widetilde{B}^{0}\Rightarrow\tau=\frac{\epsilon cl_{p}}{\widetilde{\Pi}_{0}}(\widetilde{Y}^{0}-\widetilde{B}^{0})$ (95a)
$Y^{1}(\tau)=\widetilde{Y}^{1}(\tau)=\frac{\widetilde{\Pi}_{1}}{(-\epsilon)cl_{p}}\tau+\widetilde{B}^{1}=-\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}\widetilde{Y}^{0}+\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}\widetilde{B}^{0}+\widetilde{B}^{1},$ (95b)
$Y^{I}(\tau)=\widetilde{Y}^{I}(\tau)=\frac{\widetilde{\Pi}_{I}}{(-\epsilon)cl_{p}}\tau+\widetilde{B}^{I}=-\frac{\widetilde{\Pi}_{I}}{\widetilde{\Pi}_{0}}\widetilde{Y}^{0}+\frac{\widetilde{\Pi}_{I}}{\widetilde{\Pi}_{0}}\widetilde{B}^{0}+\widetilde{B}^{I}.$ (95c)
On the other hand, $Y^{0}$ and $\widetilde{Y}^{0}$ are related by the first equation in (V.2). We can solve it for $Y^{0}$ as
$e^{4Y^{0}}$ $=\frac{\widetilde{\Pi}_{0}^{2}}{\bar{c}^{2}}\frac{e^{4\widetilde{Y}^{0}}}{(e^{4\widetilde{Y}^{0}}+\epsilon)^{2}}$ (96)
$=\begin{cases}\frac{\widetilde{\Pi}_{0}^{2}}{4\bar{c}^{2}}\frac{1}{(\mathrm{cosh}(2\widetilde{Y}^{0}))^{2}}&(\text{for }\epsilon=1)\\ \frac{\widetilde{\Pi}_{0}^{2}}{4\bar{c}^{2}}\frac{1}{(\mathrm{sinh}(2\widetilde{Y}^{0}))^{2}}&(\text{for }\epsilon=-1)\end{cases}$
Now, we introduce a new radial coordinate $R$ by
$e^{4\widetilde{Y}^{0}}=:\frac{R+r_{+}}{(-\epsilon)(R-r_{-})}$ (97)
for $R\neq r_{-}$, where $r_{\pm}$ are constants to be determined, and $\epsilon=-1$ corresponds to $R>r_{-}$ and $\epsilon=1$ to $R<r_{-}$. We want to express the metric (81) in terms of $R$. We start with evaluating the areal radius $r$:
$r^{2}$ $=\frac{l_{s}^{2}}{4}e^{4Y^{0}-4Y^{1}}$
$=\frac{l_{s}^{2}\widetilde{\Pi}_{0}^{2}}{4\bar{c}^{2}}\frac{e^{4\widetilde{Y}^{0}}}{(e^{4\widetilde{Y}^{0}}+\epsilon)^{2}}e^{4\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}\widetilde{Y}^{0}-4\left(\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}$
$=\frac{l_{s}^{2}\widetilde{\Pi}_{0}^{2}}{4\bar{c}^{2}}e^{-4\left(\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}\frac{(R+r_{+})^{1+\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}}[(-\epsilon)(R-r_{-})]^{1-\frac{\widetilde{\Pi}_{0}}{\widetilde{\Pi}_{1}}}}{(r_{+}+r_{-})^{2}}$
$=(R+r_{+})^{1+\frac{1}{\mu}}[(-\epsilon)(R-r_{-})]^{1-\frac{1}{\mu}}.$ (98)
Here, in the second line, we used the solution (95b) and the relation (96); in the third line, we applied the definition (97); and in the last line, we defined
$\mu\equiv\frac{\widetilde{\Pi}_{0}}{\widetilde{\Pi}_{1}},~r_{\pm}\equiv\frac{r_{0}}{2}(\mu\pm 1),~r_{0}\equiv\frac{l_{s}\widetilde{\Pi}_{1}}{2\bar{c}}e^{-2\left(\frac{1}{\mu}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}.$ (99)
Note that we can check $\{\mu,H_{g\phi}\}=\{r_{0},H_{g\phi}\}=0$ for (V.2), meaning that $\mu$ and $r_{0}$ are gauge invariant. Similarly, we can calculate the $tt$-component as
$4\epsilon e^{4Y^{1}}dt^{2}$ $=4e^{-4\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}\widetilde{Y}^{0}+4\left(\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}dt^{2}$
$=\epsilon\left(\frac{(-\epsilon)(R-r_{-})}{R+r_{+}}\right)^{\frac{1}{\mu}}dt_{\infty}^{2},$ (100)
where we introduced $dt_{\infty}\equiv 2e^{2\left(\frac{1}{\mu}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}dt=\frac{l_{s}\widetilde{\Pi}_{1}}{\bar{c}r_{0}}dt$. Furthermore, we can check from the definitions (97) and the trajectory (95a) and (95b) that
$d\tau$ $=\frac{cl_{p}(r_{+}+r_{-})}{4\widetilde{\Pi}_{0}}\frac{1}{(-\epsilon)(R-r_{-})(R+r_{+})}dR,$
$e^{4Y^{0}}$ $=\frac{\widetilde{\Pi}_{0}^{2}}{\bar{c}^{2}(r_{+}+r_{-})^{2}}(-\epsilon)(R-r_{-})(R+r_{+}),$
$e^{-4Y^{1}}$ $=\left(\frac{R+r_{+}}{(-\epsilon)(R-r_{-})}\right)^{\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}}e^{-4\left(\frac{\widetilde{\Pi}_{1}}{\widetilde{\Pi}_{0}}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}$ (101)
holds, enabling us to express the $\tau\tau$-component as
$-4\epsilon e^{8Y^{0}-4Y^{1}}d\tau^{2}=(-\epsilon)\left(\frac{R+r_{+}}{(-\epsilon)(R-r_{-})}\right)^{\frac{1}{\mu}}dR^{2},$ (102)
where we used the definitions (99). Thus, we reach
$ds^{2}$ $=(-\epsilon)\left(-\left[\frac{(-\epsilon)(R-r_{-})}{R+r_{+}}\right]^{\frac{1}{\mu}}dt_{\infty}^{2}\right$.
$~~~~~~~\left.+\left[\frac{R+r_{+}}{(-\epsilon)(R-r_{-})}\right]^{\frac{1}{\mu}}dR^{2}\right)+r(R)^{2}d\Omega^{2},$ (103)
where $r(R)$ is given by (V.3). Note that we have not used the Hamiltonian constraint $H_{g\phi}\approx 0$ yet.
Next, we evaluate the scalar fields $\phi_{I}=\frac{1}{\sqrt{\pi G}}\widetilde{Y}^{I}$:
$\phi_{I}(R)$ $=\frac{1}{\sqrt{\pi G}}\left(-\frac{\widetilde{\Pi}_{I}}{\widetilde{\Pi}_{0}}\widetilde{Y}^{0}+\frac{\widetilde{\Pi}_{I}}{\widetilde{\Pi}_{0}}\widetilde{B}^{0}+\widetilde{B}^{I}\right)$
$=\frac{1}{\sqrt{\pi G}}\left(\frac{\widetilde{\Pi}_{I}}{4\widetilde{\Pi}_{0}}\log\frac{(-\epsilon)(R-r_{-})}{R+r_{+}}+\frac{\widetilde{\Pi}_{I}}{\widetilde{\Pi}_{0}}\widetilde{B}^{0}+\widetilde{B}^{I}\right)$
$=\frac{\mathcal{A}_{I}}{\mu}\log\frac{(-\epsilon)(R-r_{-})}{R+r_{+}}+\phi_{I}^{(0)}.$ (104)
Here, in the first line we applied the trajectory (95c); in the second line we used the definition (97); and in the last line we defined
$\mathcal{A}_{I}\equiv\frac{1}{\sqrt{16\pi G}}\frac{\widetilde{\Pi}_{I}}{\widetilde{\Pi}_{1}},~~\phi_{I}^{(0)}\equiv\frac{1}{\sqrt{\pi G}}\left(\frac{\widetilde{\Pi}_{I}}{\widetilde{\Pi}_{0}}\widetilde{B}^{0}+\widetilde{B}^{I}\right).$ (105)
Finally, we impose the Hamiltonian constraint (V.2):
$H_{g\phi}\approx 0$ $\Rightarrow-\widetilde{\Pi}_{0}^{2}+\widetilde{\Pi}_{1}^{2}+\sum_{I}\widetilde{\Pi}_{I}^{2}\approx 0.$
$\Rightarrow\mu^{2}\equiv\left(\frac{\widetilde{\Pi}_{0}}{\widetilde{\Pi}_{1}}\right)^{2}\approx 1+16\pi G\sum_{I}\mathcal{A}_{I}^{2}.$ (106)
This provides a constraint to the physical value of $\mu$:
$\mu\geq 1,$ (107)
where $\mu=1$ holds only for $\mathcal{A}_{I}=0$.
In this solution, the energy-momentum tensor $T_{\mu\nu}=\sum_{I}(\partial_{\mu}\phi_{I}\partial_{\nu}\phi_{I}-\frac{1}{2}g_{\mu\nu}g^{\alpha\beta}\partial_{\alpha}\phi_{I}\partial_{\beta}\phi_{I})$ satisfies
$T^{t}{}_{t}=T^{R}{}_{R}=-T^{\theta}{}_{\theta}=-T^{\phi}{}_{\phi}$, (108)
with
$-T^{t}{}_{t}$ $=\frac{1}{2}g^{\tau\tau}\sum_{I}\dot{\phi}_{I}^{2}$
$=\frac{r_{0}^{2}\sum_{I}\mathcal{A}_{I}^{2}}{2(-\epsilon)}(R+r_{+})^{-2-\frac{1}{\mu}}[(-\epsilon)(R-r_{-})]^{-2+\frac{1}{\mu}},$ (109)
where we have used the configuration (V.3) and (V.3) and the $\tau$-$R$ relation (V.3).
Let us now investigate the physical properties of the solution for several cases.
(i) $\mu=1$ for $R\neq 0$: Schwarzschild metric
In this case, the constraint (V.3) means the vanishing amplitude of scalar fields: $\mathcal{A}_{I}=0$. From (V.3), then, the energy-momentum tensor becomes zero, leading to a vacuum configuration. Here, we have $r_{-}=0$ and $r_{+}=r_{0}$ (from (99)), and the relation (V.3) provides $r=R+r_{0}$ for $R\neq 0$. Thus, in the case of $\mu=1$ and $R\neq 0$, the metric (V.3) reduces to the Schwarzschild metric both for $\epsilon=-1$ and $\epsilon=+1$:
$ds^{2}=-\left(1-\frac{r_{0}}{r}\right)dt_{\infty}^{2}+\left(1-\frac{r_{0}}{r}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}.$ (110)
Here, the ADM energy $M$ is given by
$M=\frac{r_{0}}{2G}=\frac{l_{s}\widetilde{\Pi}_{1}}{4G\bar{c}}e^{-2\left(\widetilde{B}^{0}+\widetilde{B}^{1}\right)},$ (111)
where we set $\mu=1$ in the definition of $r_{0}$ (99). Although this expression is different from that of (53), we can check that $\frac{r_{0}}{2G}$ agrees with the Komar mass even in the case of $\mu\neq 1$ (see Appendix A). Note that $r_{0}$ is gauge invariant (as shown above), and so this expression of $M$ is.
(ii) $\mu>1$ and $\epsilon=-1$: a generalized JNW solution
This describes the configuration for $R>r_{-}$ given (from (V.3), (V.3), and (V.3)) by
$ds^{2}=-\left[\frac{R-r_{-}}{R+r_{+}}\right]^{\frac{1}{\mu}}dt_{\infty}^{2}+\left[\frac{R+r_{+}}{R-r_{-}}\right]^{\frac{1}{\mu}}dR^{2}+r(R)^{2}d\Omega^{2},$ (112)
with
$r(R)^{2}$ $=(R+r_{+})^{1+\frac{1}{\mu}}(R-r_{-})^{1-\frac{1}{\mu}},$
$\phi_{I}(R)$ $=\frac{\mathcal{A}_{I}}{\mu}\log\frac{R-r_{-}}{R+r_{+}}+\phi_{I}^{(0)},$ (113)
where $\mu$ satisfies the constraint (V.3). This is the JNW metric Janis:1968zz that is generalized to $n$ scalar fields.
We see the physics briefly Janis:1968zz . The energy density (V.3) is non-zero due to $\mathcal{A}_{I}\neq 0$ (from (V.3) and $\mu>1$), approaching to zero for $R\to\infty$ and to positive infinity for $R\to r_{-}$ (see Left of Fig.1).
[…]
On the other hand, the areal radius $r(R)$ (V.3) becomes large for a large $R$ and vanishes for $R\to r_{-}$ (see Right of Fig.1). Therefore, this represents an asymptotically-flat non-vacuum spacetime with a naked singularity at the center, $r=0$. It is not an isotropic physical fluid: $T^{R}_{R}\neq T^{\theta}_{\theta}=T^{\phi}_{\phi}$ and $T^{\theta}_{\theta}<0$. Note that one should be careful when taking the limit of $\mu\to 1$ at $R=r_{-}$ Janis:1968zz .
(iii) $\mu>1$ and $\epsilon=+1$: a closed universe
This case corresponds to the domain $r_{+}<R<r_{-}$, and the configuration is given by
$ds^{2}=-\left[\frac{r_{+}+R}{r_{-}-R}\right]^{\frac{1}{\mu}}dR^{2}+\left[\frac{r_{-}-R}{r_{+}+R}\right]^{\frac{1}{\mu}}dt_{\infty}^{2}+r(R)^{2}d\Omega^{2},$ (114)
with
$r(R)^{2}$ $=(r_{+}+R)^{1+\frac{1}{\mu}}(r_{-}-R)^{1-\frac{1}{\mu}},$
$\phi_{I}(R)$ $=\frac{\mathcal{A}_{I}}{\mu}\log\frac{r_{-}-R}{r_{+}+R}+\phi_{I}^{(0)}.$ (115)
We examine the meaning of this configuration. In the metric (114), $R$ is timelike and $t_{\infty}$ spacelike, and it should be natural to interpret $T^{R}{}_{R}$ as the energy-density $\rho$ and $T^{t}_{t}$ as the radial pressure. From (108), then $\rho=p$ holds, where $p=T^{t}{}_{t}=T^{\theta}{}_{\theta}=T^{\phi}{}_{\phi}$. Therefore, this is a causal-limit isotropic fluid Zeldovich:1961sbr . Next, we see from (V.3) that $r(R)^{2}$ vanishes at $r=-r_{+}$ and $r=r_{-}$ as shown in Left of Fig.2. We also check from (108) and (V.3) that the energy-density $\rho=-T^{R}{}_{R}|_{\epsilon=1}$ diverges at $r=-r_{+}$ and $r=r_{-}$ as shown in Right of Fig.2.
[…]
The spacetime volume is finite:
$\mathcal{V}_{\rm 4D}$ $=\int d^{4}x\sqrt{-g}$
$=4\pi\Delta t_{\infty}\int^{r_{-}}_{-r_{+}}dR~r(R)^{2}$
$=\frac{\pi^{2}l_{p}^{6}}{6l_{0}^{2}}\widetilde{\Pi}_{1}^{3}e^{-4\left(\frac{1}{\mu}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}\frac{\mu^{2}-1}{\sin\frac{\pi}{\mu}},$ (116)
where $\Delta t_{\infty}=\frac{l_{0}l_{s}\widetilde{\Pi}_{1}}{\bar{c}r_{0}}=\frac{l_{p}^{2}\widetilde{\Pi}_{1}}{r_{0}}$ and the definitions (99) have been used. Note that it is still non-zero after taking $\mu\to 1$, $\lim_{\mu\to 1}\mathcal{V}_{\rm 4D}=\frac{\pi l_{p}^{6}}{3l_{0}^{2}}\widetilde{\Pi}_{1}^{3}e^{-4\left(\frac{1}{\mu}\widetilde{B}^{0}+\widetilde{B}^{1}\right)}$, but the meaning is unclear yet since the finiteness comes from the singular behavior for $\mu\to 1$ in the last factor of (V.3).
Thus, this configuration represents a Kantowski-Sachs-type closed universe for a causal-limit isotropic fluid that has initial and final singularities at $R=-r_{+},~r_{-}$ corresponding to a big bang and big crunch, respectively. Indeed, setting $n=1$ and using the scalar field $\phi_{I=1}$ as a “clock” (instead of $R$), the metric (114) agrees with the one obtained in Ref.Chiou:2008eg . The new point we find is that the closed universe and the JNW spacetime can be unified into the single expression (V.3); for the domain $r_{+}<R<r_{-}$, it represents the closed universe while for $r_{-}<R<\infty$ the JNW spacetime. In this sense, the closed universe corresponds to the “interior” of the JNW metric, although the two spacetimes are disconnected at $R=r_{-}$ where the singularity exists in both sides.
Let us conclude this section by summarizing the results. For a coupled system of gravity and $n$ massless scalar fields, we have employed the different choice of canonical variables and gauge from the ones used in the previous cases, and established the $(2+n)$-D-Schrödinger symmetry. In the vacuum case, this leads to a partial evidence for the covariance of the symmetry in the mini-superspace. Then, we have obtained the generalized JNW solution and the closed universe solution as its “interior”. They are characterized by the mass parameter $r_{0}$ and the parameter $\mu$ related to the amplitude of the scalar fields. In terms of the canonical variables, the solutions are determined by the values of the generators of the symmetry: $(\widetilde{\Pi}_{0},\widetilde{\Pi}_{1},\widetilde{B}^{0},\widetilde{B}^{1})$. Therefore, the symmetry transformation can change the values. We will discuss this point in the next section.
VI Interpretation of the Symmetries under Hamiltonian Constraints
In the previous sections, we have identified Schrödinger symmetry in the mini-superspace models. However, these systems differ from non-relativistic free particle systems due to the presence of Hamiltonian constraints. Indeed, they significantly complicate the interpretation of the symmetry.
We first summarize what we have done so far. For the mini-superspace models with matter fields, we have employed the canonical-transformation method and re-expressed the Hamiltonian constraints $H\approx 0$ (for certain choices of $N$) as
$H=\frac{1}{2}\mathcal{G}^{ab}P_{a}P_{b}+\lambda\equiv H_{0}+\lambda\approx 0,$ (117)
to reach Schrödinger symmetries. Here, $\lambda$ is a constant characterizing the dynamics, and $P_{a}$ is used to denote $\widetilde{P}_{a}$ or $\widetilde{\Pi}_{a}$ for simplicity.
Now, before studying the physical meaning of this symmetry, we here list the points we should keep in mind when considering systems with gauge constraints in general. Consider a symmetry generator $G$ satisfying the condition (6). Then, we have
  • (I)
    When a state satisfying the constraints is transformed by $G$, does the transformed state satisfy the constraints again?
  • (II)
    Does applying a transformation $G$ to a quantity $X$ after using the constraints yield the same result as using the constraints after transforming $X$ by $G$?
  • (III)
    States connected by gauge transformations are physically equivalent. When transformed by $G$, are they still physically equivalent?
Let us add a little more. For (I), the physical states are restricted on a hypersurface in the phase space determined by the constraints, and some transformation $G$ may move such states outside the surface, rendering them no longer physical ones. For (II), consider, for example, transforming the ADM mass $M$ by $G=\widetilde{B}^{-}$. The definition (53) contains $\widetilde{P}_{-}$, and the mass changes by the transformation. However, if we use the constraint (43) and express $\widetilde{P}_{-}$ in $\widetilde{P}_{+}$ first, then the mass does not. For (III), physically equivalent states are represented by a single gauge orbit. If a transformation $G$ moves a gauge orbit to another one, that transformation is gauge-independent. In other words, the transformation of a gauge-invariant observable $\mathcal{O}$ by such a $G$ is gauge-invariant, where $\mathcal{O}$ commutes with the gauge constraints.
The points (I)–(III) are satisfied if a symmetry generator $G$ commutes with the gauge constraints, as shown below in an example. Actually, the transformation induced by such a generator does not alter the theory; that is, it acts as a symmetry transformation that leaves the Hamiltonian invariant.
On the other hand, for symmetry generators $G$ that do not commute with the gauge constraints, the three points require careful attention. In our system (117), there exist such symmetry generators: $B^{a}$, $Q_{0}$, and $Q_{-}$. One possible interpretation of transformations by $Q_{0}$ and $Q_{-}$ involves incorporating the transformation of the lapse function $N$ Achour2019 . In contrast, for $B^{a}$, to the best of our knowledge, no such interpretation exists.
We will propose another possible interpretation: such non-commuting symmetries transform the theory itself, and the transformed configurations are physical solutions of this new theory. First, we will develop a general argument for this view (Sec. VI.1), and then explicitly demonstrate it with examples (Sec. VI.2 and Sec.VI.3).
VI.1 Symmetry generating a new theory
We argue that the symmetries by non-commuting generators $G$ with the Hamiltonian constraint $H$ transform the theory.
We start with reviewing Noether’s theorem without considering $H\approx 0$. We first check the role of the symmetry condition (6). Under an infinitesimal canonical transformation generated by a function $G(x,p,t)$, the transformed canonical variables $(x^{\prime},p^{\prime})$ and the transformed Hamiltonian $H^{\prime}(x^{\prime},p^{\prime},t)$ are given by
$x^{\prime i}=x^{i}+\epsilon\{x^{i},G\},$ (118a)
$p^{\prime}_{i}=p_{i}+\epsilon\{p_{i},G\},$ (118b)
$H^{\prime}(x^{\prime},p^{\prime},t)=H(x,p)+\epsilon\frac{\partial G}{\partial t},$ (118c)
where $H(x,p)$ is the original Hamiltonian and $\epsilon$ is a small parameter. If $G$ satisfies the symmetry condition (6), it follows from (118) that
$H^{\prime}(x,p)=H(x,p).$ (119)
That is, the Hamiltonian preserves its functional form. Therefore, symmetry transformations leave the theory invariant. In particular, symmetry transformations can be interpreted as transformations of the initial conditions of the solution trajectory.
We next see this explicitly for finite canonical transformations of $X^{a}$ by a generator $G$ of the Schrödinger symmetry in the Hamiltonian (117). Noting $X^{a}=B^{a}+\mathcal{G}^{ab}P_{b}\tau$ from (5), it can be calculate as
$X^{\prime a}$ $=e^{s\{\cdot,G\}}\triangleright X^{a}$
$=e^{s\{\cdot,G\}}\triangleright B^{a}+\mathcal{G}^{ab}(e^{s\{\cdot,G\}}\triangleright P_{b})\tau,$ (120)
where a canonical transformation of a function $f$ by a generator $G$ for a finite parameter $s$ is denoted as
$e^{s\{\cdot,G\}}\triangleright f:=f+s\{f,G\}+\frac{1}{2!}s^{2}\{\{f,G\},G\}$
$+\frac{1}{3!}s^{3}\{\{\{f,G\},G\},G\}+\cdots.$ (121)
Here, $e^{s\{\cdot,G\}}\triangleright B^{a}$ and $e^{s\{\cdot,G\}}\triangleright P_{b}$ are expressed only in terms of the generators of the Schrödinger symmetry, since the symmetry algebra is closed. Note also that the generators can be built from $P_{a}$ and $B^{a}$ (see Sec.II.1). Now, for a solution to the canonical equation, we have a uniform linear motion in the configuration space $\{X^{a}\}$:
$X^{a}(\tau)=\beta^{a}+\mathcal{G}^{ab}\alpha_{b}\tau,$ (122)
which corresponds to replacing the generators $(P_{a},B^{a})$ with integration constants $(\alpha_{a},\beta^{a})$ as in Sec.II.1. Then, the transformed trajectory is given from (VI.1) by
$X^{\prime}(\tau)$ $=(e^{s\{\cdot,G\}}\triangleright B^{a})|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}}$
$~~~~~~~~~+\mathcal{G}^{ab}(e^{s\{\cdot,G\}}\triangleright P_{b})|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}}\tau$
$\equiv\beta^{\prime a}+\mathcal{G}^{ab}\alpha^{\prime}_{a}\tau$ (123)
which is another uniform linear motion. Thus, a uniform linear motion (a solution to the canonical equation) $X^{a}(\tau)$ is mapped to another one $X^{\prime a}(\tau)$ by the symmetry transformations. This is the meaning of the dynamical symmetry.
Now, let us consider the effect of $H\approx 0$. Suppose that the original trajectory $X^{a}(\tau)$ (122) satisfies it:
$H|_{P_{a}\to\alpha_{a}}=\frac{1}{2}\mathcal{G}^{ab}\alpha_{a}\alpha_{b}+\lambda=0.$ (124)
Then, $X^{a}(\tau)$ presents a physical solution of the theory with the Hamiltonian (117). Noting that $\mathcal{G}^{ab}$ is Lorentzian, the constraint (124) can be interpreted as a “dispersion relation” between “energy” $\alpha_{0}$ and “momentum” $\alpha_{\rm other}$ for the characteristic “mass” $\lambda$.
Due to this constraint, all the components of $\alpha_{a}$ are not independent for physical states. On the other hand, under the gauge transformation generated by the Hamiltonian (117), $\alpha_{a}$ remains invariant, while $\beta^{a}$ is gauge-dependent and carries unphysical degrees of freedom. This clear separation into $\alpha_{a}$ and $\beta^{a}$ is a key advantage of working with the Hamiltonian of the form (117).
Next, let us examine whether the transformed trajectory $X^{\prime a}(\tau)$ (VI.1) satisfies the Hamiltonian constraint, $H|_{P_{a}\to\alpha_{a}^{\prime}}=0$, or not. First, we can check, by explicit calculation,
$e^{s\{\cdot,G\}}\triangleright H_{0}=\frac{1}{2}\mathcal{G}^{ab}(e^{s\{\cdot,G\}}\triangleright P_{a})(e^{s\{\cdot,G\}}\triangleright P_{b}).$ (125)
Noting $(e^{s\{\cdot,G\}}\triangleright P_{a})|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}}\equiv\alpha^{\prime}_{a}$, therefore, in order to see $H|_{P_{a}\to\alpha_{a}^{\prime}}=0$, we just have to show $(e^{s\{\cdot,G\}}\triangleright H)|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}}=0$. Here, the expression
$e^{s\{\cdot,G\}}\triangleright H=\lambda+H_{0}+s\{H_{0},G\}+\frac{1}{2}s^{2}\{\{H_{0},G\},G\}+\cdots.$ (126)
means from (125) that
$\{H_{0},G\}=0~{\rm and}~H|_{P_{a}\to\alpha_{a}}=0\Longrightarrow H|_{P_{a}\to\alpha^{\prime}_{a}}=0.$ (127)
For this case, therefore, the transformed solution $X^{\prime}(\tau)$ is a new physical solution within the original theory. (We will see this explicitly in Sec.VI.2.) This is the reason why point (I) in the last subsection is satisfied for symmetry generators that commute with $H$.
Similarly, when $\{G,H_{0}\}=0$, points (II) and (III) hold. For point (II), since $G$ doest not change $H$, even if we rewrite a quantity $X$ using $H=0$ into another expression before applying the transformation $G$, the result will match that obtained by applying $G$ to the original expression of $X$. For point (III), $G$ and $H$ commute, and so do $e^{s\{\cdot,G\}}\triangleright$ and $e^{s^{\prime}\{\cdot,H\}}\triangleright$, meaning that a gauge orbit is mapped to anothoer one by $e^{s\{\cdot,G\}}\triangleright$.
However, if $\{H_{0},G\}\neq 0$, the transformed trajectory $X^{\prime}(\tau)$ does not satisfy $H|_{P_{a}\to\alpha^{\prime}_{a}}=0$ and thus does not become a physical solution to the original theory with $H$. To interpret such a transformation as a kind of symmetry, we consider the following. For a physical solution with $(\alpha_{a},\beta^{a})$ to the original theory with $H$ (that is, $H|_{P_{a}\to\alpha_{a}}=0$), we use a generator $G$ of the Schrödinger symmetry and introduce a new characteristic constant as [^5] Here, the value of $s$ is chosen in a specific manner. See below for examples.
$\lambda_{\rm New}:=\lambda+(H_{0}-e^{s\{\cdot,G\}}\triangleright H_{0})|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}}$ (128)
to define a new Hamiltonian or new dispersion relation:
$H_{\rm New}:=H_{0}+\lambda_{\rm New}.$ (129)
Then, for a trajectory $X^{\prime}(\tau)$ generated by $G$, we can use the property (125) and calculate
$H_{\rm New}|_{P_{a}\to\alpha^{\prime}_{a},B^{a}\to\beta^{\prime a}}$ $=\lambda_{\rm New}+(e^{s\{\cdot,G\}}\triangleright H_{0})|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}}$
$=\lambda+H_{0}|_{P_{a}\to\alpha_{a}}$
$=0.$ (130)
This shows that the transformed trajectory $X^{\prime}(\tau)$ is a physical solution to the newnew theory $H_{\rm New}$. [^6] A theory is specified by a Hamiltonian constraint $H\approx 0$ and canonical-transformation relations $\widetilde{X}^{a}(X,P),\widetilde{P}_{a}(X,P)$, which enable us to translate the uniform linear motion in $\{\widetilde{X}^{a}\}$ into the spacetime metric $g_{\mu\nu}$ in terms of $X^{a}$. Here, the new theory has a different constraint but the same relations. Note that the new Hamiltonian also holds the Schrödinger symmetry, since $\lambda_{\rm new}$ is constant and the analysis we have done so far is applicable to $H_{\rm New}$. Thus, this procedure generates a new theory preserving the Schrödinger symmetry and transforms a physical solution of the original theory to a physical solution of the new theory. (We will demonstrate this explicitly in Sec.VI.3.)
From a perspective of this symmetry transformation that changes $\alpha_{a}$, $\beta^{a}$, and $\lambda$ simultaneously, let us consider points (I), (II) and (III). Point (I) is achieved by adopting the idea of moving the constraint surface itself while preserving the way of transformation of a point in the phase space. For Point (II), $H=H_{0}+\lambda$ remains invariant according to this generalized transformation, and no ambiguity appears in transforming a quantity $X$.
We finally discuss point (III). We here have to be careful with transformation of not only a gauge-invariant observable $\mathcal{O}$ but also $\lambda$, since $\lambda$ transforms as in Eq.(128) and determines the theory. From Eq. (128) and the fact that $H$ does not involve $B^{a}$, the transformation of $\lambda$ induced by generators that depend at most on linear forms of $B^{a}$ is gauge independent. Actually, for a function $F(P)$ and a generator $G$,
$\{F(P),G\}|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}}=\left(-\frac{\partial F}{\partial P_{a}}\frac{\partial G}{\partial B^{a}}\right)|_{P_{a}\to\alpha_{a},B^{a}\to\beta^{a}}$ (131)
is independent of $\beta^{a}$ if $\frac{\partial G}{\partial B^{a}}$ does not include $B^{a}$, and inductively, terms of any order in $s$ in (128) become independent of $\beta^{a}$. Note that one of the Schrödinger symmetry generators, $Q_{-}$, contains a quadratic term in $B^{a}$, and the resulting new theory is gauge-dependent. [^7] The presence of $\beta^{a}$ does not necessarily imply gauge dependence. There exist physical quantities that involve $\beta^{a}$ but are nonetheless gauge-independent. For example, $J_{ab}$.
On the other hand, the transformation of gauge-invariant observables $\mathcal{O}$ can be gauge dependent. For instance, the transformation of $\mathcal{O}=J_{ab}$ by $B^{a}$ is gauge dependent, since their commutation relation is linear on $B^{a}$. Once we determine the value of $\beta^{a}$, the transformation is uniquely determined. There is room for discussion regarding the role of the gauge in this view of symmetry.
Thus, the above perspective can be considered as one interpretation of Noether’s theorem in a theory with the Hamiltonian constraint $H\approx 0$ and symmetry generators $G$ that do not commute with $H$. It is worth noting that transformations commuting with $H$ can be regarded, from this perspective, as a special case where $\lambda$ remains unchanged. Furthermore, in this interpretation, all symmetry transformations satisfying the condition (6) preserve the property (119) of leaving the time evolution operator $H_{0}$ invariant for certain choices of $N$. Therefore, this interpretation of the $\lambda$ transformation can be viewed as an extension of the symmetry perspective—where initial conditions are changed without altering the time-evolution operator $H_{0}$—to the transformation of $\lambda$.
VI.2 $\{G,H\}=0$: Solution transformations
We use the mini-superspace models discussed so far and demonstrate that generators $G$ with $\{G,H\}=0$ map a physical solution to another in the same theory with $H$, where both satisfy $H\approx 0$. Among the generators of the Schrödinger symmetry, $\widetilde{P}_{a}$ and $\widetilde{J}_{ab}$ are of this type. We will pick up particularly interesting cases.
Ex.1: $\widetilde{J}_{-A}$ in the Maxwell-field mini-superspace
We consider the Maxwell-field mini-superspace with the Hamiltonian constrain (69),
$H_{gE}$ $=-\frac{2}{\epsilon cl_{p}}\widetilde{P}_{+}\widetilde{P}_{-}+\frac{1}{2cl_{p}}\widetilde{P}_{A}^{2}-\frac{cl_{p}}{2l_{s}^{2}},$
and study the transformation of the (A)dS Reissner-Nordström solution (IV.2) by $\widetilde{J}_{-A}$. In particular, let us map this to the (A)dS Schwarzschild solution (III.3) with a different value of mass $M$.
We start by using the 3D Schrödinger algebra (72) and calculating finite canonical transformation of interest by $\widetilde{J}_{-A}=\widetilde{\mathcal{G}}_{-+}^{E}\widetilde{B}^{+}\widetilde{P}_{A}-\widetilde{\mathcal{G}}_{AA}^{E}\widetilde{B}^{A}\widetilde{P}_{-}=\frac{(-\epsilon)cl_{p}}{2}\widetilde{B}^{+}\widetilde{P}_{A}-cl_{p}\widetilde{B}^{A}\widetilde{P}_{-}$:
$\widetilde{P}^{\prime}_{+}$ $=e^{s\{\cdot,\widetilde{J}_{-A}\}}\triangleright\widetilde{P}_{+}=\widetilde{P}_{+}+\frac{\epsilon cl_{p}s}{2}\widetilde{P}_{A}+\frac{\epsilon c^{2}l_{p}^{2}s^{2}}{4}\widetilde{P}_{-}$
$\widetilde{P}^{\prime}_{-}$ $=e^{s\{\cdot,\widetilde{J}_{-A}\}}\triangleright\widetilde{P}_{-}=\widetilde{P}_{-}$
$\widetilde{P}^{\prime}_{A}$ $=e^{s\{\cdot,\widetilde{J}_{-A}\}}\triangleright\widetilde{P}_{A}=\widetilde{P}_{A}+cl_{p}s\widetilde{P}_{-}$
$\widetilde{B}^{\prime+}$ $=e^{s\{\cdot,\widetilde{J}_{-A}\}}\triangleright\widetilde{B}^{+}=\widetilde{B}^{+}$
$\widetilde{B}^{\prime-}$ $=e^{s\{\cdot,\widetilde{J}_{-A}\}}\triangleright\widetilde{B}^{-}=\widetilde{B}^{-}-cl_{p}s\widetilde{B}^{A}+\frac{\epsilon c^{2}l_{p}^{2}s^{2}}{4}\widetilde{B}^{+}$
$\widetilde{B}^{\prime A}$ $=e^{s\{\cdot,\widetilde{J}_{-A}\}}\triangleright\widetilde{B}^{A}=\widetilde{B}^{A}-\frac{\epsilon cl_{p}s}{2}\widetilde{B}^{+}$. (132)
Suppose that the original (A)dS-RN solution is characterized by a set of the integration constants $(\alpha_{+},\alpha_{-},\alpha_{A};\beta^{+},\beta^{-},\beta^{A})$ satisfying the Hamiltonian constraint (69): $H_{gE}|_{\widetilde{P}_{a}\to\alpha_{a}}=0$.
Now, we set the transformation parameter $s$ such that the electric charge (75) vanishes: $\mathcal{Q}^{\prime}=\sqrt{\frac{\pi}{G}}\frac{l_{s}}{\bar{c}}\alpha^{\prime}_{A}=0$, where $\alpha^{\prime}_{A}=P^{\prime}_{A}|_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=0$ with
$s=-\frac{1}{cl_{p}}\frac{\alpha_{A}}{\alpha_{-}}\equiv s_{*}.$ (133)
Then, the other charges transform from (VI.2) as follows:
$\alpha^{\prime}_{+}$ $=\widetilde{P}^{\prime}_{+}|_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=\alpha_{+}+\frac{(-\epsilon)\alpha_{A}^{2}}{4\alpha_{-}}$
$\alpha^{\prime}_{-}$ $=\widetilde{P}^{\prime}_{-}|_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=\alpha_{-}$
$\beta^{\prime+}$ $=\widetilde{B}^{\prime+}|_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=\beta^{+}$
$\beta^{\prime-}$ $=\widetilde{B}^{\prime-}|_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=\beta^{-}+\frac{\alpha_{A}}{\alpha_{-}}\beta^{A}+\frac{\epsilon\alpha_{A}^{2}}{4\alpha_{-}^{2}}\beta^{+}$
$\beta^{\prime A}$ $=\widetilde{B}^{\prime A}|_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=\beta^{A}+\frac{\epsilon\alpha_{A}}{2\alpha_{-}}\beta^{+}.$ (134)
These change the value of mass $M$ (53): $M^{\prime}=\frac{\epsilon l_{s}}{G\bar{c}^{2}}\alpha^{\prime}_{+}(\alpha^{\prime}_{+}\beta^{\prime+}-\alpha_{-}^{\prime}\beta^{\prime-})$. On the other hand, the new charges still satisfy the original Hamiltonian constraint:
$H_{gE}|_{\widetilde{P}_{a}\to\alpha^{\prime}_{a}}$ $=-\frac{2}{\epsilon cl_{p}}\alpha^{\prime}_{+}\alpha^{\prime}_{-}+\frac{1}{2cl_{p}}\alpha_{A}^{{}^{\prime}2}-\frac{cl_{p}}{2l_{s}^{2}}$
$=-\frac{2}{\epsilon cl_{p}}\left(\alpha_{+}+\frac{(-\epsilon)\alpha_{A}^{2}}{4\alpha_{-}}\right)\alpha_{+}-\frac{cl_{p}}{2l_{s}^{2}}$
$=H_{gE}|_{\widetilde{P}_{a}\to\alpha_{a}}=0.$ (135)
Therefore, we reach the AdS-Schwarzschild solution with mass $M^{\prime}$. This shows that the symmetry transformation by $J_{A-}$ for $s=s_{*}$ maps the AdS-RN solution with $(M,\mathcal{Q})$ to the AdS-Schwarzschild solution with $M^{\prime}$ in the same Hamiltonian constraint $H_{gE}\approx 0$.
Ex.2: $\widetilde{J}_{1\phi}$ in the scalar-field mini-superspace model.
We next provide an example in the scalar-field mini-superspace model (V.2). For simplicity, we consider the case of $n=1$ here:
$H_{g\phi}$ $=\frac{1}{2(-\epsilon)cl_{p}}\left(-\widetilde{\Pi}_{0}^{2}+\widetilde{\Pi}_{1}^{2}+\widetilde{\Pi}_{\phi}^{2}\right).$ (136)
Let us apply the symmetry transformation by $\widetilde{J}_{1\phi}$ and map the JNW solution (and its interior) to the Schwarzschild solution.
First, the finite canonical transformation of the relevant generators is given by, through the Schrödinger algebra associated with (93),
$\widetilde{\Pi}^{\prime}_{1}$ $=e^{s\{\cdot,\widetilde{J}_{1\phi}\}}\triangleright\widetilde{\Pi}_{1}=\widetilde{\Pi}_{1}\cos{\Theta}-\widetilde{\Pi}_{\phi}\sin{\Theta}$
$\widetilde{\Pi}^{\prime}_{\phi}$ $=e^{s\{\cdot,\widetilde{J}_{1\phi}\}}\triangleright\widetilde{\Pi}_{\phi}=\widetilde{\Pi}_{\phi}\cos{\Theta}+\widetilde{\Pi}_{1}\sin{\Theta}$
$\widetilde{B}^{\prime 1}$ $=e^{s\{\cdot,\widetilde{J}_{1\phi}\}}\triangleright\widetilde{B}^{1}=\widetilde{B}^{1}\cos{\Theta}+\widetilde{B}^{\phi}\sin{\Theta}$
$\widetilde{B}^{\prime\phi}$ $=e^{s\{\cdot,\widetilde{J}_{1\phi}\}}\triangleright\widetilde{B}^{\phi}=\widetilde{B}^{\phi}\cos{\Theta}-\widetilde{B}^{1}\sin{\Theta},$ (137)
where $\Theta\equiv(-\epsilon)cl_{p}s$.
We assume that the original solution metric (V.3) with the condition (V.3) is specified by a set of the constants $(\alpha_{0},\alpha_{1},\alpha_{\phi};\beta^{0},\beta^{1},\beta^{\phi})$, where the Hamiltonian constraint (136) holds: $H_{g\phi}|_{\widetilde{\Pi}_{a}\to\alpha_{a}}=0$.
We then choose the parameter $s=\frac{\Theta}{(-\epsilon)cl_{p}}$ such that $\alpha^{\prime}_{\phi}=\widetilde{\Pi}^{\prime}_{\phi}|_{\widetilde{\Pi}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=0$. That is,
$\Theta_{*}=(-\epsilon)cl_{p}s_{*}=\arctan\left({-\frac{\alpha_{\phi}}{\alpha_{1}}}\right).$ (138)
Applying this to $\alpha^{\prime}_{a}=\widetilde{\Pi}^{\prime}_{a}|_{\widetilde{\Pi}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}$ and using $\cos^{2}{\Theta_{*}}=\frac{1}{1+\alpha_{\phi}^{2}/\alpha_{1}^{2}}$, we can calculate
$-\alpha^{\prime 2}_{0}+\alpha^{\prime 2}_{1}+\alpha^{\prime 2}_{\phi}$
$=$ $-\alpha_{0}^{2}+\alpha_{1}^{2}\cos^{2}{\Theta_{*}}+\alpha_{\phi}^{2}\sin^{2}{\Theta_{*}}-2\alpha_{1}\alpha_{\phi}\cos{\Theta_{*}}\sin{\Theta_{*}}$
$=$ $\cos^{2}{\Theta_{*}}\left[-\frac{\alpha_{0}^{2}}{\cos^{2}{\Theta_{*}}}+\alpha_{1}^{2}+\alpha_{\phi}^{2}\tan^{2}{\Theta_{*}}-2\alpha_{1}\alpha_{\phi}\tan{\Theta_{*}}\right]$
$=$ $\cos^{2}{\Theta_{*}}\left[-\alpha_{0}^{2}\left(1+\frac{\alpha_{\phi}^{2}}{\alpha_{1}^{2}}\right)+\alpha_{1}^{2}+\alpha_{\phi}^{2}\frac{\alpha_{\phi}^{2}}{\alpha_{1}^{2}}-2\alpha_{1}\alpha_{\phi}\frac{-\alpha_{\phi}}{\alpha_{1}}\right]$
$=$ $\cos^{2}{\Theta_{*}}\left[-\alpha_{0}^{2}+\alpha_{1}^{2}+\alpha_{\phi}^{2}+\left(\frac{\alpha_{\phi}}{\alpha_{1}}\right)^{2}(-\alpha_{0}^{2}+\alpha_{\phi}^{2})+\alpha_{\phi}^{2}\right]$
$=$ $0,$ (139)
where in the last step we have used $H_{g\phi}|_{\widetilde{\Pi}_{a}\to\alpha_{a}}=0$. Therefore, the transformed configuration $Y^{\prime a}(\tau)=\beta^{\prime a}+\mathcal{G}^{ab}_{\phi}\alpha^{\prime}_{b}\tau$ satisfies the original Hamiltonian constraint. Here, the $\mu$ parameter (V.3) changes as $\mu=\frac{\alpha_{0}}{\alpha_{1}}>1\to\frac{\alpha^{\prime}_{0}}{\alpha^{\prime}_{1}}=1$ because of $\alpha^{\prime}_{\phi}=0$. Therefore, this is the Schwarzschild solution with a different value of mass (111), $M^{\prime}=\frac{l_{s}\alpha^{\prime}_{1}}{4G\bar{c}}e^{-2\left(\beta^{\prime 0}+\beta^{\prime 1}\right)}$.
VI.3 $\{G,H\}\neq 0$: Theory transformations
Finally, we present two examples where a theory itself changes under transformations by a symmetry generator $G$ that does not commute with the Hamiltonian constraint $H$. $\widetilde{B}^{a}$, $\widetilde{Q}_{0}$, and $\widetilde{Q}_{-}$ belong to this type. We then discuss what kind of theory it is.
Ex.3: $\widetilde{B}^{+}$ in the Maxwell-field mini-superspace model.
We discuss the symmetry transformation by the Galilei boost $\widetilde{B}^{+}$ in the Maxwell-field mini-superspace model (69), where the constant $\lambda$ in (117) corresponds to $\frac{-cl_{p}}{2l_{s}^{2}}$. First, suppose again that the original AdS-RN solution is specified by a set $(\alpha_{+},\alpha_{-},\alpha_{A};\beta^{+},\beta^{-},\beta^{A})$ that satisfies $H_{gE}|_{\widetilde{P}_{a}\to\alpha_{a}}=0$. We then define the new parameter (128) by
$\lambda_{\rm New}:$ $=\frac{-cl_{p}}{2l_{s}^{2}}+(H_{0}-e^{s\{\cdot,\widetilde{B}^{+}\}}\triangleright H_{0})|_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}$
$=\frac{-cl_{p}}{2l_{s}^{2}}+\frac{2\alpha_{-}}{(-\epsilon)cl_{p}}s_{*}$
$=-\frac{cl_{p}}{2l_{s}^{2}}(1+k_{*}),$ (140)
for a certain value $s_{*}$, where we have employed the 3D Schrödinger algebra (72) and defined $k_{*}\equiv\frac{4\epsilon\alpha_{-}}{\bar{c}^{2}}s_{*}$. This gives the new dispersion relation:
$H^{\rm New}_{gE}=\frac{2}{(-\epsilon)cl_{p}}\widetilde{P}_{+}\widetilde{P}_{-}+\frac{1}{2cl_{p}}\widetilde{P}_{A}^{2}-\frac{cl_{p}}{2l_{s}^{2}}(1+k_{*}).$ (141)
At the same time, the configuration $X^{a}(\tau)=\beta^{a}+\mathcal{G}^{ab}_{E}\alpha_{b}\tau$ is transformed to the new one $X^{\prime a}(\tau)=\beta^{\prime a}+\mathcal{G}^{ab}_{E}\alpha^{\prime}_{b}\tau$, where
$\alpha^{\prime}_{+}$ $=\left(e^{s\{\cdot,\widetilde{B}^{+}\}\}}\triangleright\widetilde{P}_{+}\right)_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=\alpha_{+}-s*$
$\alpha^{\prime}_{-}$ $=\left(e^{s\{\cdot,\widetilde{B}^{+}\}\}}\triangleright\widetilde{P}_{-}\right)_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=\alpha_{-}$
$\alpha^{\prime}_{A}$ $=\left(e^{s\{\cdot,\widetilde{B}^{+}\}\}}\triangleright\widetilde{P}_{A}\right)_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=\alpha_{A}$
$\beta^{\prime a}$ $=\left(e^{s\{\cdot,\widetilde{B}^{+}\}\}}\triangleright\widetilde{B}^{a}\right)_{\widetilde{P}_{a}\to\alpha_{a},\widetilde{B}^{a}\to\beta^{a},s\to s_{*}}=\beta^{a}.$ (142)
Then, $\alpha^{\prime}_{a}$ satisfy not the original Hamiltonian constraint but the new one:
$H^{\rm New}_{gE}|_{\widetilde{P}_{a}\to\alpha^{\prime}_{a}}$
$=\frac{2}{(-\epsilon)cl_{p}}(\alpha_{+}-s*)\alpha_{-}+\frac{1}{2cl_{p}}\alpha_{A}^{2}-\frac{cl_{p}}{2l_{s}^{2}}+\frac{2\alpha_{-}}{(-\epsilon)cl_{p}}s_{*}$
$=0,$ (143)
because of $H_{gE}|_{\widetilde{P}_{a}\to\alpha_{a}}=0$. Therefore, this symmetry transformation generates the new theory with $H_{gE}^{\rm New}$ and maps the solution of the original theory $H_{gE}$ into another solution of the new one $H_{gE}^{\rm New}$.
Now, two questions arise: (i) What kind of spacetime is the obtained configuration? and (ii) What is the new theory $H_{gE}^{\rm New}$?
We first note from footnote 6 that the original relations $\widetilde{X}^{a}(X,P),\widetilde{P}_{a}(X,P)$ are kept in the new theory, and the same geometric interpretation holds except for the parts involving the Hamiltonian constraint. Then, the calculation for $g_{rr}$ (IV.2) is modified as
$-\epsilon N^{2}d\tau^{2}$ $=\frac{1}{\frac{1}{\bar{c}^{2}}\left(-\frac{4}{\epsilon}\alpha^{\prime}_{+}\alpha^{\prime}_{-}+\alpha^{\prime 2}_{A}\right)-\frac{2GM^{\prime}}{r}+\frac{G\mathcal{Q}^{2}}{4\pi r^{2}}-\frac{\Lambda}{3}r^{2}}dr^{2}$
$\approx\frac{1}{1+k_{*}-\frac{2GM^{\prime}}{r}+\frac{G\mathcal{Q}^{2}}{4\pi r^{2}}-\frac{\Lambda}{3}r^{2}}dr^{2}.$ (144)
Here, $H^{\rm New}_{gE}|_{\widetilde{P}_{a}\to\alpha^{\prime}_{a}}\approx 0$ is applied, and the value of the mass (53) changes from $M$ to $M^{\prime}=\frac{\epsilon l_{s}}{G\bar{c}^{2}}\alpha^{\prime}_{+}(\alpha^{\prime}_{+}\beta^{+}-\alpha_{-}\beta^{-})$ while the electric charge $\mathcal{Q}$ doesn’t. We can also have a similar modification for $g_{tt}$ (see (IV.2)). Therefore, we reach a different metric from the AdS-RN metric:
$ds^{2}\approx$ $-\left(1+k_{*}-\frac{2GM^{\prime}}{r}+\frac{G\mathcal{Q}^{2}}{4\pi r^{2}}-\frac{\Lambda}{3}r^{2}\right)dt_{\infty}^{2}$
$+\left(1+k_{*}-\frac{2GM^{\prime}}{r}+\frac{G\mathcal{Q}^{2}}{4\pi r^{2}}-\frac{\Lambda}{3}r^{2}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}.$ (145)
Even for the case of $\Lambda=0$, it is not asymptotically flat. A new term due to $k_{*}\neq 0$ appears in the energy density and radial pressure:
$G^{t}{}_{t}$ $=G^{r}{}_{r}=-\Lambda-\frac{G\mathcal{Q}^{2}}{4\pi r^{4}}+\frac{k_{*}}{r^{2}},$
$G^{\theta}{}_{\theta}$ $=G^{\phi}{}_{\phi}=-\Lambda+\frac{G\mathcal{Q}^{2}}{4\pi r^{4}},$
$R$ $=4\Lambda-\frac{2k_{*}}{r^{2}}.$ (146)
This is a solution to the new theory $H^{\rm New}_{gE}$.
Next, to identify what the new theory is, let us review how we have obtained it. We first note from (VI.3) that the shift $\frac{cl_{p}}{2l_{s}^{2}}\to-\frac{cl_{p}}{2l_{s}^{2}}(1+k_{*})$ corresponds to generating the new theory (141). We then remember that the term $\frac{cl_{p}}{2l_{s}^{2}}$ in the original theory (69) comes from the second term, $\frac{cl_{p}}{l_{s}^{2}}$, in the reduced gravity Langrangian (28). Therefore, we can see that a 2D covariant term
$\Delta S^{\rm New}_{gE}$ $=\frac{1}{16\pi G\hbar}\int d^{4}x\sqrt{-g}\frac{8\pi k_{*}}{\mathcal{A}}$
$=\int d^{2}x\sqrt{-g_{2d}}\frac{k_{*}}{2l_{p}^{2}}$ (147)
makes the shift effect. Here, $\mathcal{A}=4\pi r^{2}$, and $\sqrt{-g_{2d}}$ is the area element for the 2D normal part to the spatial sphere. Thus, the new metric (VI.3) solves the equation of motion for the Einstein-gravity coupled with Maxwell field plus the 4D cosmological constant $\Lambda$ and the 2D cosmological constant $\Lambda_{2d}\equiv-\frac{k_{*}}{4l_{p}^{2}}$. Although more studies on this new theory would be needed, at least, it keeps the 3D Schrödinger symmetry (as seen at the end of Sec.VI.1).This theory-symmetry method would be one procedure for generating a modified gravity theory with Schrödinger symmetry.
Ex.4: $\widetilde{B}^{\phi}$ in the scalar-field mini-superspace model.
We now consider the transformation by $\widetilde{B}^{\phi}$ in the single scalar-field mini-superspace model (136), and discuss the new theory and configuration.
In particular, for the original solution metric (V.3) with constants $(\alpha_{0},\alpha_{1},\alpha_{\phi};\beta^{0},\beta^{1},\beta^{\phi})$ satisfying the Hamiltonian constraint $H_{g\phi}\approx 0$, we study the finite transformation by $\widetilde{B}^{\phi}$ for $s=\alpha_{\phi}\equiv s_{*}$. For this parameter we evaluate $\lambda_{\rm New}$ (128) and find
$H_{g\phi}^{\rm New}=\frac{1}{2(-\epsilon)cl_{p}}\left(-\widetilde{\Pi}_{0}^{2}+\widetilde{\Pi}_{1}^{2}+\widetilde{\Pi}_{\phi}^{2}\right)+\frac{1}{2(-\epsilon)cl_{p}}\alpha_{\phi}^{2}$ (148)
Because of $\lambda_{\rm New}={\rm const.}\neq 0$, this theory is not Einstein gravity plus a scalar field but keeps the 3D Schrödinger symmetry.
By a similar calculation to (VI.3), we can check that $\alpha^{\prime}_{0}=\alpha_{0}$, $\alpha^{\prime}_{1}=\alpha_{1}$, and $\alpha^{\prime}_{\phi}=0$, meaning that the scalar field $\phi$ is not excited now. Nevertheless, this new configuration satisfies the new Hamiltonian constraint:
$H_{g\phi}^{\rm New}|_{\widetilde{\Pi}_{a}\to\alpha^{\prime}_{a}}$ $=\frac{1}{2(-\epsilon)cl_{p}}\left(-\alpha^{\prime 2}_{0}+\alpha^{\prime 2}_{1}+\alpha^{\prime 2}_{\phi}+\alpha_{\phi}^{2}\right)$
$=\frac{1}{2(-\epsilon)cl_{p}}\left(-\alpha^{2}_{0}+\alpha^{2}_{1}+0+\alpha^{2}_{\phi}\right)=0.$ (149)
Therefore, the new configuration is given by the JNW-type metric (V.3) with $\mu^{2}=\left(\frac{\alpha^{\prime}_{0}}{\alpha^{\prime}_{1}}\right)^{2}\approx 1+\left(\frac{\alpha_{\phi}}{\alpha_{1}}\right)^{2}$, where the new constraint (VI.3) is applied. It does not excite the scalar field $\phi$ but has the same metric as the original configuration.
What is this new theory? One simple interpretation is to consider it as the Einstein gravity coupled with the scalar field $\phi$ in the background of another scalar field $\psi$ with the same excitation $\alpha_{\phi}$ as the original scalar-field configuration. More precisely, considering the $Y^{\phi}$ solution (95c) and the relation $\widetilde{Y}^{\phi}=Y^{\phi}=\sqrt{\frac{\pi}{\hbar}}l_{p}\phi$, we set the background field as
$\psi\to\psi_{\rm back}(\tau)=\sqrt{\frac{\hbar}{\pi}}\frac{\alpha_{\phi}}{(-\epsilon)cl_{p}^{2}}\tau+\psi_{0},$ (150)
where $\psi_{0}$ is an arbitrary constant. Then, we add
$S_{\psi}^{\rm(back)}[g_{\mu\nu}]=-\frac{1}{2\hbar}\int d^{4}x\sqrt{-g}g^{\mu\nu}\partial_{\mu}\psi_{\rm back}(\tau)\partial_{\nu}\psi_{\rm back}(\tau)$ (151)
to the Einstein-Hilbert action coupled with single massless scalar field $\phi_{I=1}=\phi$ (V.1) where $\Lambda=0$, $m_{I}=0$, and $I=1$. This is the new theory. In other words, this can be regarded as an effective theory obtained by introducing another field $\psi$ with the same excitation as the excitation of field $\phi$ in the original configuration and then integrating it out. The new configuration with the JNW metric discussed above is a solution of this new theory, due to the background field $\psi_{\rm back}$ which replaces the effect of the original field $\phi$. Better understanding the physical meaning of this would require further research.
VII Conclusions and Prospects
Symmetry should be a useful concept in exploring the principles of quantum gravity. The gauge symmetry of spacetime is general covariance, and Hamiltonian constraints play a central role in the dynamics of gravity. Then, what is the physical symmetry in the configuration space of gravity? Hamiltonian constraints define an infinite-dimensional configuration space, “superspace”. Imposing geometric symmetries (such as homogeneity) on spacetime yields a finite-dimensional configuration space, “mini-superspace”. This reduction procedure can be viewed as a kind of fluid limit for (quantum) gravitational configurations. Generically, when many microscopic degrees of freedom are coarse-grained into a fluid description, symmetries like volume-preserving diffeomorphism invariance often emerge, independent of the details of the constituents. Therefore, it is expected that some common symmetry may emerge in various mini-superspaces, potentially carrying the universal symmetry of quantum gravity. Indeed, 2D Schrödinger symmetry was obtained in the homogeneous isotropic universe model and the spherically symmetric static vacuum model. In this paper, to investigate its robustness, we analyzed spherically symmetric static mini-superspace models coupled with matter fields at a classical level. Due to spherical symmetry and staticity, these become one-dimensional dynamical systems with Hamiltonian constraint $H=N\mathcal{H}\approx 0$, where the radial coordinate is the evolution parameter.
To find the symmetries of such systems, we developed the method utilizing canonical transformations. This method not only applies to a broader class of models than conventional approaches but also possesses the advantage of being systematically treated within the standard framework of canonical theory. First, applying this method to the spherically symmetric static mini-superspace model with Maxwell field and cosmological constant, we demonstrated that 3D Schrödinger symmetry emerges for a certain choice of $N$ and the $X$ coordinate system in mini-superspace. Here, the dimension of the mini-superspace increases by one for Maxwell field $A_{0}$, and the solution is given by the AdS-RN solution. Next, we analyzed the spherically symmetric static mini-superspace model coupled to $n$ massless scalar fields. Using a different $N$ and coordinate system $Y$ than before, we showed that (n+2)(n+2)D Schrödinger symmetry arises. Solving this model yielded the JNW solution and its “interior” solution in a unified manner. Taking the matter decoupling limit, both become vacuum mini-superspace models, yielding the 2D Schrödinger symmetry for different lapses and mini-superspace coordinate systems. This suggests the covariance of Schrödinger symmetry in the mini-superspace. Thus, the result that Schrödinger symmetry is obtained for mini-superspaces with different dimensions and different types of matter fields, along with its covariance, supports the robustness of the emergence of Schrödinger symmetry in the fluid limit.
We finally explored the physical meaning of Schrödinger symmetry under the Hamiltonian constraint $H\approx 0$. Generators $G$ commuting with $H$ transform a solution satisfying $H\approx 0$ to another one (dynamical symmetry). On the other hand, this interpretation does not hold for generators $G$ that do not commute with $H$. Then, we proposed one possible interpretation: such a symmetry transforms the theory itself, and the configuration obtained by the transformation is a solution of the new theory. We explained this with examples. While this interpretation remains open to debate, it provides a first step toward understanding the meaning of symmetry under Hamiltonian constraints.
Let us now consider several future prospects.
First, it is crucial to investigate the robustness of Schrödinger symmetry in a broader context. Our current discussion is limited to cases reducible to the free-particle Hamiltonian without potential via canonical transformations. However, potentials of special forms, such as $V(x)=\frac{k}{x^{2}}$, also preserve Schrödinger symmetry Boyer:1974qt ; Nikitin:2017gtu ; Duval:2024eod . Therefore, it is important to consider such more general cases and investigate whether other matter fields, such as massive scalar field or Dirac field, possess Schrödinger symmetry, or reconsider whether the scalar-field model in terms of the $X$ coordinate in footnote 4 leads to the symmetry. Conversely, it would be interesting to assume such special potentials and examine the spacetime structure derived from them or consider the corresponding modified gravity theories.
A related problem is the exploration for dynamics that eliminates singularities. One approach is to construct an effective theory inspired by DeWitt’s idea DeWitt:1967yk . One would examine what behavior trajectories on the mini-superspace should exhibit to avoid singularities, and search for potentials that can realize it while preserving Schrödinger symmetry. The corresponding theory should contain effective dynamics for resolving singularities.
Also, a deeper understanding of the interpretation of symmetries that generate theories is needed. For this, it should be useful to consider an extended phase space formulation where time (and its conjugate, energy) is incorporated as canonical variables Fahn:2018ahm ; struckmeier2005hamiltonian . This could allow the symmetry condition (6) to be written in the form of an extended Poisson bracket, $\{G,H\}_{*}=0$, enabling a more systematic investigation of the symmetries in systems with $H=0$, including generators that do not commute with $H$.
Furthermore, there are various possibilities for quantizing the mini-superspace models using Schrödinger symmetry. One approach is to study quantization techniques specific to particular potentials Essin:2006sic and to consider how they apply to the models. Another method adopts the idea of group quantization Isham:1988bx to quantize the Schrödinger group. Here, the dimension of our mini-superspace models is larger than 2, and it would be more non-trivial to construct the Casimirs and representations than the previous study Sartini:2021ktb .
In particular, quantization of the scalar-field mini-superspace model is interesting. The interior region of the horizon in the Schwarzschild metric corresponds—via swapping two coordinates $t$ and rr— to a Kantowski-Sachs-type cosmological spacetime, where no matter field exists, of course. From this corresponding, quantization models of black holes are expected to exhibit a strong link with quantum cosmology models Modesto:2005zm . On the other hand, a scalar field can be used as a “clock” in relational formalism to quantize models Domagala:2010bm ; Ashtekar:2011ni . One of our interesting results in the scalar-field mini-superspace model is that, the JNW solution and the Kantowski-Sachs-type spacetime (with the causal-limit fluid) are unified into the single metric (V.3), meaning that the latter corresponds to the “interior” of the former. In particular, both possess the Schrödinger symmetry. Therefore, the quantization of this model based on the symmetry and this interior-exterior relationship should provide a promising avenue for connecting quantum black hole models and quantum universe models.
[…]
Finally, it is crucial to examine the Schrödinger symmetry found in our mini-superspace models from the perspective of a hydrodynamic approximation derived from a full theory of quantum gravity. Ref.Oriti:2024elx focuses on the Schrödinger symmetry in the FLRW mini-superspace model and discusses the potential for removing cosmological singularities and realizing accelerated expansion through corrections to the quantum fluid equations that preserve the symmetry. It would be interesting to check whether a similar mechanism works in the spherically-symmetric static mini-superspace models. This test could serve as a crucial benchmark for assessing the universality of the hydrodynamic picture in quantum gravity. On the other hand, Ref.BenAchour:2023dgj pointed out the existence of Schrödinger symmetry in the spherically-symmetric static vacuum mini-superspace and suggested that nonlinear terms preserving the symmetry could serve as quantum corrections. However, the form of such corrections may depend on the choice of the lapse function. Indeed, the symmetry they identified corresponds to only one of the two possible choices of the lapse function and mini-superspace coordinate we found. It would be intriguing to consider different lapses and coordinates and investigate whether there are differences in the physical effects arising from the correction terms. Furthermore, we showed that Schrödinger symmetry persists in spherically-symmetric static spacetimes coupled with matter fields. If the emergence of Schrödinger symmetry in such a mini-superspace is rigorously linked to the validity of the hydrodynamic approximation in a full quantum gravity theory incorporating matter fields, the findings of this study would provide significant implications.
Acknowledgments
We would like to thank Jibril Ben Achour, Che-Yu Chen, Kristina Giesel, Philipp Höhn, Christy Kelly, Etera Livine, Francesco Sartini, and Shinji Tsujikawa for fruitful and inspiring discussions. T.S. is supported by Waseda University Special Research Projects (Nos. 2025C-493) and the 15th Early Bird Program of Waseda University. Y.Y. appreciated the wonderful hospitality of the members of the Institute for Theoretical Physics III at the Friedrich-Alexander-University Erlangen-Nuremberg, where he had a fruitful discussion and completed this work. Y.Y. was partially supported by Japan Society for the Promotion of Science (No.21K13929).
Appendix A Mass and Electric Charge
We check the consistency for the formulae of mass $M$, (53) and(111), and electric charge $\mathcal{Q}$ (75).
Mass
The Komar mass formula can evaluate the ADM energy of an asymptotically-flat stationary spacetime in a covariant way Wald:1984rg :
$M_{\rm Komar}\equiv-\frac{1}{8\pi G}\int_{S_{\infty}}dS_{\mu\nu}\nabla^{\mu}t^{\nu},$ (152)
where $t^{\mu}$ is the timelike Killing vector satisfying $t^{2}\xrightarrow[r\to\infty]{}-1$, and $dS_{\mu\nu}$ is the area element for a 2-sphere $S_{\infty}$ in an asymptotically flat vacuum region. For a spherically-symmetric static spacetime with the metric (26), we can calculate it as
$M_{\rm Komar}$ $=-\frac{1}{8\pi G}\int_{S_{\infty}}\nabla^{t_{\infty}}t^{\tau}dS_{t_{\infty}\tau}\times 2$
$=-\frac{1}{4\pi G}\int_{S_{\infty}}g^{t_{\infty}t_{\infty}}\nabla_{t_{\infty}}t^{\tau}\sqrt{-g}d\theta d\phi$
$=\frac{1}{8\pi G}4\pi r^{2}\frac{1}{\sqrt{-g_{2d}}}\partial_{\tau}(-g_{t_{\infty}t_{\infty}}),$ (153)
where $g_{2d}$ is the determinant of the 2D metric normal to 2-sphere, and $dt_{\infty}=kdt$ is the time coordinate at $r=\infty$ with a constant $k$.
First, we prove that the mass (53) for $\Lambda=0$ and $\epsilon=-1$ agrees with the Komar formula (153). We start by expressing the metric (26) in terms of $t_{\infty}$:
$ds^{2}=\epsilon\frac{4X^{+}(\tau)}{X^{-}(\tau)}\frac{\widetilde{P}_{+}^{2}}{\bar{c}^{2}}dt^{2}_{\infty}-\epsilon\frac{X^{-}(\tau)}{4X^{+}(\tau)}d\tau^{2}+\frac{l_{s}^{2}}{4}{X^{-}(\tau)}^{2}d\Omega^{2},$ (154)
where we used the gauge $N=\sqrt{\frac{X^{-}}{4X^{+}}}$ and $dt_{\infty}\equiv\frac{\bar{c}}{\widetilde{P}_{+}}dt$. We can then calculate the Komar mass (153) as
$M_{\rm Komar}$ $=\frac{\epsilon l_{s}^{2}\widetilde{P}_{+}}{2G\bar{c}}\left(X^{+}\dot{X}^{-}-X^{-}\dot{X}^{+}\right)$
$=\frac{l_{s}}{G\bar{c}^{2}}P_{+}(B^{-}P_{-}-B^{+}P_{+}).$ (155)
Here, noting that $(\widetilde{X}^{a},\widetilde{P}_{a})=(X^{a},P_{a})$ holds in the case of $\Lambda=0$, we have applied the canonical equations (III.3): $\dot{X}^{\pm}=-\frac{2}{\epsilon cl_{p}}P_{\mp}={\rm const.}$ and $X^{\pm}=-\frac{2}{\epsilon cl_{p}}P_{\mp}\tau+B^{\pm}$. This agrees with the definition of $M$ (53) for $\Lambda=0$ and $\epsilon=-1$.
Next, we show that in the Y-coordinate, the Komar formula (153) becomes the simpler formula (111). In terms of $dt_{\infty}=\frac{l_{s}\widetilde{\Pi}_{1}}{\bar{c}r_{0}}dt\equiv\frac{1}{k}dt$ and the gauge (85), the Y-metric (81) is expressed as
$ds^{2}=4\epsilon\left(-e^{8Y^{0}-4Y^{1}}d\tau^{2}+k^{2}e^{4Y^{1}}dt_{\infty}^{2}\right)+\frac{l_{s}^{2}}{4}e^{4Y^{0}-4Y^{1}}d\Omega^{2}.$ (156)
The formula (153) can be evaluated as
$M_{\rm Komar}$ $=\frac{(-\epsilon)}{2G}l_{s}^{2}k\dot{Y}^{1}$
$=\frac{(-\epsilon)}{2G}l_{s}^{2}\frac{\bar{c}r_{0}}{l_{s}\widetilde{\Pi}_{1}}\frac{\widetilde{\Pi}_{1}}{(-\epsilon)cl_{p}}$
$=\frac{r_{0}}{2G},$ (157)
where we used the definition of $k$ and the trajectory (95b). This is consistent with the definition (111).
Electric charge
The standard electric charge is given by
$\mathcal{Q}_{\rm{elecric}}\equiv\int_{\Sigma}d\Sigma_{\mu}j^{\mu}=\frac{1}{2}\int_{S}dS_{\mu\nu}F^{\mu\nu}.$ (158)
Here, $j^{\mu}$ is electric current, $d\Sigma_{\mu}$ is the volume element of a hypersurface $\Sigma$, and we applied Maxwell’s equations $\nabla_{\nu}F^{\mu\nu}=j^{\mu}$ and Stokes’ theorem.
In the metric (154), we can evaluate this as
$\mathcal{Q}_{\rm{electric}}$ $=\frac{1}{2}\int_{S}dS_{t_{\infty}\tau}F^{t_{\infty}\tau}\times 2$
$=\int_{S}d\theta d\phi\sqrt{-g}g^{t_{\infty}t_{\infty}}g^{\tau\tau}F_{t_{\infty}\tau}$
$=4\pi r^{2}\frac{1}{\sqrt{-g_{2d}}}\dot{A}_{t_{\infty}}$
$=\sqrt{\frac{\pi}{G}}l_{s}^{2}(X^{-})^{2}\dot{X}^{A},$ (159)
where we recalled $X^{A}\equiv\sqrt{\pi G}A_{t}$. Using the relation (IV.1) of $\widetilde{X}^{A}$ and $X^{A}$ and the canonical equation, we have
$\dot{X}^{A}$ $=\dot{\widetilde{X}}^{A}+\frac{\epsilon\widetilde{P}_{A}}{2\widetilde{P}_{+}}\dot{\widetilde{X}}^{-}\left(1-\frac{1}{(X^{-})^{2}}\right)$
$=\frac{\widetilde{P}_{A}}{cl_{p}}+\frac{\epsilon\widetilde{P}_{A}}{2\widetilde{P}_{+}}\frac{-2}{\epsilon cl_{p}}\widetilde{P}_{+}\left(1-\frac{1}{(X^{-})^{2}}\right)$
$=\frac{1}{cl_{p}^{2}(X^{-})^{2}}\widetilde{P}_{A}.$ (160)
Substituting this into (A), we obtain
$\mathcal{Q}_{\rm{electric}}=\sqrt{\frac{\pi}{G}}\frac{l_{s}}{\bar{c}}\widetilde{P}_{A},$ (161)
which matches the definition of $\mathcal{Q}$ (75).